### v1

# Importance Guided Data Augmentation for Neural-Based Code Understanding

Abstract
Pre-trained code models lead the era of code intelligence. Many models have been designed with impressive performance recently. However, one important problem, data augmentation for code data that automatically helps developers prepare training data lacks study in the field of code learning. In this paper, we introduce a general data augmentation framework, GenCode, to enhance the training of code understanding models. GenCode follows a generation-and-selection paradigm to prepare useful training codes. Specifically, it uses code transformation techniques to generate new code candidates first and then selects important ones as the training data by importance metrics. To evaluate the effectiveness of GenCode with a general importance metric – loss value, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5). Compared to the state-of-the-art (SOTA) code augmentation method, MixCode, GenCode produces code models with 2.92% higher accuracy and 4.90% robustness on average.
1 Introduction
Deep learning (DL) for code has been one of the successful DL applications over the years. The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection Zhou et al. (2019), is commonly known as code learning. From the model level, many DL models specifically designed for code learning have been proposed, such as CodeBERT Feng et al. (2020), GraphCodeBERT Guo et al. (2020), and CodeT5 Wang et al. (2021). From the business perspective, multiple related products such as Copilot  [^1] https://github.com/features/copilot and Tabnine  [^2] https://www.tabnine.com have been widely used to help our daily development. As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  Lu et al. (2021) and program classification Puri et al. (2021), DL already achieved significantly better results than traditional SE techniques.
However, most of the existing research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models Sun et al. (2022). The challenge behind this is preparing training data for code models is heavy work. Although today developers can easily collect raw code data from the open-source community such as GitHub [^3] https://github.com, data labeling is still expensive, especially for code-related tasks. It needs human effort with professional domain knowledge.
Techniques can be directly used for code augmentation, such as code refactoring methods and text transformation methods. Unfortunately, existing studies Bielik and Vechev (2020); Yu et al. (2022) stated that code refactoring methods have limited advantages in boosting the training of code models. Besides,  Dong et al. (2023b) revealed that almost all the data augmentation methods cannot produce code models with good robustness. Moreover, a recent work, MixCode Dong et al. (2023a), that follows the spirit of Mixup Zhang et al. (2018) to do code augmentation achieved promising results, but it can be only used for classification tasks. Therefore, there are no effective and reliable code augmentation methods that can be generalized to different tasks.
To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision Cubuk et al. (2020); Yang et al. (2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding. Specifically, at each training epoch, GenCode first generates training candidates by using different code transformation methods. Here, we consider different types of code transformation methods such as semantic-preserving code refactoring methods and syntax-breaking text transformation methods. After that, GenCode selects valuable training data from the candidates with the guidance of importance metrics. The important metric used in GenCode is simply defined as the loss value of code data produced by the code model. Finally, the selected samples are used for the model training. We conducted comprehensive experiments to study how GenCode can improve the accuracy and robustness of code models. The experimental results demonstrated that, compared to the SOTA method, MixCode, GenCode can produce code models with a 2.92% higher accuracy and 4.90% better robustness on average.
To summarize, the main contributions of this paper are:
  • We propose the first generation-and-selection-based framework, GenCode, which automatically enhances the training data and identifies the most important data from the augmented set for boosting the code understanding.
  • We empirically demonstrate that GenCode can produce more accurate and robust code understanding models compared to existing data augmentation methods and conduct an ablation study to explore how different importance scores affect the effectiveness of GenCode.
2 Background and Related Work
2.1 Deep Learning for Code
Since source code can be seen as natural language or structural graph Allamanis et al. (2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated bug detection Dinella et al. (2020), automated program classification Puri et al. (2021), and code clone detection Wang et al. (2020). The key problem in deep learning for code is how to represent raw code to machine-readable inputs. Researchers designed different methods for code representation that can be roughly divided into two groups, sequence-based representation and graph-based representation.
Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens. For example, given a code snippet, “def func(x, y)”, sequence-based representation transforms it to a token list as “[’def’, ’func’, ’(’, ’x’, ’y’, ’)’]”. After that, each token in the token list will be mapped to a pre-defined index value, e.g., [1, 25, 34, 5100, 640, 33]. Finally, these index values will be fed into the code-embedding models for the learning.
Graph-based representation learns the structural information of code. Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information Dinella et al. (2020); Yasunaga and Liang (2020). For example, GraphCodeBERT combines the sequence-based representation and data flow information of code as a whole to train code models with better performance.
In this work, we focus on both sequence-based and graph-based representation and study the powerful pre-trained code models CodeBERT, GraphCodeBERT, and CodeT5 that achieved impressive results on code understanding tasks.
2.2 Data Augmentation for Code
Data augmentation is a well-studied technique in the machine learning field to increase the size and diversity of training data by only modifying the existing labeled data, thus, without further human effort. For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym Feng et al. (2021).
Recently, data augmentation gained attention from software engineering researchers. Various data augmentation methods designed for source code have been proposed Gao et al. (2020); Yefet et al. (2020); Allamanis et al. (2021); Pour et al. (2021); Bui et al. (2021); Wang et al. (2022); Park et al. (2023); Yang et al. (2023a), with the most general framework involving the use of code refactoring, e.g., local variable renaming, duplication, and dead store, to create more code data without breaking the semantics of the original codes. Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information Lacerda et al. (2020).
Existing works Bielik and Vechev (2020); Yu et al. (2022); de Paula Rodrigues et al. (2023); Dong et al. (2023b, 2024) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models. For instance, a recent empirical study Dong et al. (2024) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation. Besides, the SOTA code data augmentation method, MixCode which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring. However, as a variant of Mixup specifically designed for code data, MixCode is limited to applications in code classification tasks, as it necessitates the label with one-hot encoding.
Compared to existing works, we propose a simple yet effective code augmentation framework that is not only unaffected by the dataset and model constraints but also applicable to both classification and non-classification tasks.
3 Definition
We first define the problem of data augmentation for code models.
Definition 3.1 (Program Code [^4] For simplicity, we use code in the paper. ).
A code is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: $content$ contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
Definition 3.2 (Code Model).
A code model is defined as a 3-tuple $M=\left(parameter,\mathcal{P},Y\right)$, where $parameter$ contains all the weights in code model $M$, $\mathcal{P}$ is the input code, and $Y$ is the label of $\mathcal{P}$ given a specific downstream code task.
Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 Wang et al. (2021).
Definition 3.3 (Code Augmentation).
Given a code $P$, a code transformation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)$.
Here, code transformation methods can be any method that can change the content of program statements such as text data transformation methods (e.g., synonym replacement and code refactoring methods (e.g., dead if statement adding).
Definition 3.4 (Semantic-preserving/breaking Code Data Augmentation).
Given a code model $M$, a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, semantic-preserving (breaking) code augmentation indicates $\mathcal{P}.se=\mathcal{P^{\prime}}.se$ $M(\mathcal{P})=M(\mathcal{P^{\prime}})$ ($\mathcal{P}.se\neq\mathcal{P^{\prime}}.se$ $M(\mathcal{P})\neq M(\mathcal{P^{\prime}})$).
Definition 3.5 (Syntax-preserving/breaking Code Augmentation).
Given a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, syntax-preserving (breaking) code augmentation indicates $\mathcal{P}.sy=\mathcal{P^{\prime}}.sy$ ($\mathcal{P}.sy\neq\mathcal{P^{\prime}}.sy$).
4 Assumption and Preliminary Study
GenCode selects the most valuable data for the code model training. This approach is grounded on the assumption that the data that code models have less confidence in is more useful for model training. To check whether this assumption is valid or not, we conducted a preliminary study to explore if there is a correlation between the accuracy of the trained code model and the confidence of the model on the training data. Specifically, the study is conducted on the Google Code Jam (GCJ) Yang et al. (2022) dataset and CodeBERT pre-trained code model. The detailed steps are as follows:
Step 1. We initialize the code model and use code augmentation methods to generate additional data as the training candidate set. The code augmentation methods will be introduced in Section 5.2.
Step 2. We randomly select a number of data from the training candidate set. Here the selected number is the same as the size of the original training data. We repeat this selection 100 times and collect 100 groups of training data. After that, for each group of training data, we compute and record their average loss values.
Step 3. We train the code model using each group of training data respectively for a few epochs and thus save 100 trained code models. We then test each code model using the original test data and save 100 accuracies.
Step 4. We use the Pearson Correlation Coefficient (PCC) Schober et al. (2018) metric to check the correlation between the loss values and the accuracies.
Table 1: Correlation between loss values and code model accuracy.
Coefficient P-value
Initialized (epoch 0) 0.6075 7.9687e-12
Middle (epoch 10) 0.2440 0.0125
Well-trained (epoch 20) 0.1657 0.0926
Table 1 summarizes the results of our preliminary study. Here, we considered three types of code models, the initialed model, the model that has been trained for a few epochs but not converged, and the well-trained code model, to check if our assumption can stand in different scenarios. The results indicate that there is a positive correlation between the loss values of data and the accuracy of code models trained by using these data. From the initialized model to the well-trained model, this correlation becomes weaker but still exists. Therefore, we can propose an effective and reliable data augmentation framework to boost the training of code models based on this phenomenon.
Finding: There is a positive correlation between the loss values of code data and the accuracy of code models trained by using these data. When the code model is randomly initialized, this correlation is significant (with a P-value less than 0.01).
5 Methodology
5.1 Overview
The goal of GenCode is to produce code models with better performance on test data by training models on carefully generated data. Given a training set of program $\mathcal{P}_{train}$, its transformed version $\mathcal{P^{\prime}}_{train}$, and a set of test programs $\mathcal{P}_{test}$, after the training process $\mathsf{train}\left(\cdot\right)$ , $\varrho(M^{\prime}(\mathcal{P}_{test}))>\varrho(M(\mathcal{P}_{test}))$, where $M.parameter=\mathsf{train}(M,\mathcal{P}_{train})$ $M^{\prime}.parameter=\mathsf{train}(M^{\prime},\mathcal{P^{\prime}}_{train})$, and $\varrho\left(\cdot\right)$ is the performance measurement.
GenCode follows the generation-and-selection workflow and utilizes different types of code transformation methods to generate diverse code data first, then selects code data that are most useful for training as the training data. Algorithm 1 shows how GenCode works. In each training epoch, GenCode automatically generates more code data (Line 7 to Line 9, Definition 3.3) and then selects the most important ones (Line 10 to Line 13) to train the code model (Line 14). Here, the data importance is measured by the loss values produced by the code model which is the most common indicator to check how the model performs on the data. Overall, GenCode contains two main steps, utilizing code data augmentation methods to compose the search space and importance-guided data selection.
Algorithm 1 GenCode
1:$M$: initialized DNN model
2:$\mathcal{P},Y$: original code data and labels
3:$\mathcal{T}=\left\{t\right\}:$ a set of code augmentation methods
4:$M$: trained model
5:for $run\leftarrow\{0,\ldots,\#epochs\}$ do
6:     $\mathcal{P^{\prime}},Y^{\prime}\ \leftarrow\phi,\phi$triangleright initialization
7:     for $t\in\mathcal{T}$ do
8:         $\mathcal{P^{\prime}}\leftarrow\mathcal{P^{\prime}}\cup t\left(\mathcal{P}\right)$
9:         $Y^{\prime}\leftarrow Y^{\prime}\cup Y$triangleright code augmentation
10:     $I\leftarrow\mathsf{Importance}\left(M,\mathcal{P^{\prime}}\right)$triangleright computing importance
11:     $Index\leftarrow\mathsf{argSort}\left(I\right)\left[:||\mathcal{P}||\right]$triangleright sort by importance
12:     $X_{train}\leftarrow\mathcal{P^{\prime}}\left[Indext\right]$
13:     $Y_{train}\leftarrow Y^{\prime}\left[Indext\right]$
14:     $M\leftarrow\mathsf{Train}\left(M,X_{train},Y_{train}\right)$
15:return $M$
5.2 Search Space
Even though the code has strict syntax constraints, recent work Dong et al. (2023b) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training. To enhance the diversity of source code within the feature space, therefore, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 3.4 and Definition 3.5).
GenCode adapts code refactoring methods as semantic/syntax-preserving methods. Code refactoring methods follow the constraints of programming languages to edit the code while keeping the main functionality of the code. On the other hand, text augmentation methods are used as syntax-breaking methods in GenCode which are widely used in the NLP field. For example, text augmentation methods can randomly delete some tokens in the code. However, even though the syntax has been broken, the code model can still successfully predict those codes since it reads the code as a sequence. In total, as shown in Table 2, GenCode currently supports 18 code refactoring methods and 5 text augmentation methods. Besides, Figure 1 provides an example of the training candidate that is to generate the search space. In the syntax-preserving, we present a code refactoring method named API Renaming, which involves modifying the function name from ”main()” to ”even()”. In the syntax-breaking, we demonstrate another instance of API Renaming, where two different statements within a program are randomly selected and their positions swapped. As depicted in this example, the third and fourth lines have been interchanged. Appendix B detailed introduces the specifics of these atomic data augmentation operators.
Table 2: Atomic data augmentation operators adopted in GenCode.
Syntax-preserving Syntax-breaking
API renaming
Filed enhancement
Synonym replacement Random insertion Random swap Random deletion Back-translation
Arguments adding
For loop enhancement
Arguments renaming
If enhancement
Dead for adding
Local variable adding
Dead if adding
Local variable renaming
Dead if-else adding
Method name renaming
Dead switch adding
Plus zero
Dead while adding
Print adding
Duplication Return optimal
Figure 1: An example of atomic data augmentation operators.
Table 3: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data and ASR $\downarrow$ (%) on test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
GCJ Refactory Java250 BigCloneBench
Model DA method Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$
No Aug 90.96 ± 0.07 33.31 56.13 96.21 ± 0.13 36.31 42.31 96.36 ± 0.09 40.55 53.37 96.91 ± 0.12 9.74 24.61
Refactor 91.74 ± 0.11 35.56 52.84 95.95 ± 0.19 32.64 40.89 96.44 ± 0.12 38.35 52.56 96.97 ± 0.21 8.55 18.29
SR 70.66 ± 0.11 35.46 52.71 96.71 ± 0.11 39.25 53.91 96.21 ± 0.02 42.01 53.79 97.02 ± 0.15 12.52 26.11
RI 59.91 ± 0.21 40.78 65.76 96.43 ± 0.18 38.45 45.57 96.15 ± 0.06 43.15 55.07 97.38 ± 0.23 8.31 30.08
RS 93.24 ± 0.13 28.26 52.44 96.73 ± 0.15 38.63 45.98 96.49 ± 0.04 39.36 50.79 96.92 ± 0.12 13.41 30.13
RD 76.98 ± 0.09 35.65 55.46 96.46 ± 0.18 30.23 41.32 96.61 ± 0.12 38.31 52.42 97.11 ± 0.19 9.58 30.15
BT 82.74 ± 0.15 40.38 55.95 94.31 ± 0.38 62.98 69.87 96.06 ± 0.06 42.21 53.32 97.13 ± 0.17 11.29 26.32
MixCode 92.21 ± 0.25 29.28 52.53 96.76 ± 0.31 31.04 41.01 96.62 ± 0.23 36.14 50.67 - - -
CodeBERT GenCode 96.23 ± 0.18 22.33 46.62 99.32 ± 0.14 23.75 31.25 98.89 ± 0.09 29.89 40.65 98.93 ± 0.11 7.12 15.83
No Aug 93.96 ± 0.15 31.23 56.02 96.84 ± 0.16 22.97 30.26 96.41 ± 0.11 23.16 42.66 96.88 ± 0.09 4.19 12.22
Refactor 91.72 ± 0.21 30.31 56.58 95.49 ± 0.21 24.26 32.57 96.53 ± 0.15 23.13 42.42 97.11 ± 0.23 6.05 6.17
SR 89.46 ± 0.13 40.02 66.68 95.57 ± 0.13 28.39 47.15 96.46 ± 0.14 22.92 42.18 97.14 ± 0.16 6.13 10.16
RI 80.47 ± 0.21 43.51 65.75 94.51 ± 0.25 25.08 32.87 96.52 ± 0.09 22.37 42.33 97.35 ± 0.17 2.23 6.28
RS 94.76 ± 0.15 30.18 54.75 97.88 ± 0.15 21.85 29.67 96.44 ± 0.14 22.68 41.98 97.54 ± 0.12 6.03 11.94
RD 91.71 ± 0.26 31.73 60.92 96.53 ± 0.19 20.35 26.98 96.56 ± 0.21 22.39 41.84 97.16 ± 0.21 6.18 8.27
BT 80.47 ± 0.05 41.65 62.99 95.99 ± 0.27 35.56 46.23 96.39 ± 0.13 22.91 42.32 97.11 ± 0.16 8.18 14.12
MixCode 94.19 ± 0.25 30.26 54.49 97.29 ± 0.15 14.89 26.53 96.89 ± 0.19 20.03 42.01 - - -
GraphCodeBERT GenCode 97.93 ± 0.14 26.43 48.87 99.45 ± 0.17 15.78 21.92 99.16 ± 0.13 18.76 33.81 99.23 ± 0.08 1.86 4.62
No Aug 95.02 ± 0.07 30.78 55.67 97.22 ± 0.17 22.45 29.69 96.98 ± 0.08 31.76 49.82 97.38 ± 0.13 10.32 22.97
Refactor 92.21 ± 0.14 30.12 55.93 96.43 ± 0.22 28.76 35.45 96.92 ± 0.11 30.66 47.12 97.24 ± 0.12 9.32 22.05
SR 90.14 ± 0.21 39.21 51.84 94.67 ± 0.14 28.24 36.26 96.53 ± 0.05 28.65 46.01 97.31 ± 0.08 15.87 25.98
RI 78.77 ± 0.05 41.02 63.46 95.02 ± 0.15 24.78 33.61 96.11 ± 0.09 27.14 46.76 97.46 ± 0.13 10.43 29.76
RS 95.66 ± 0.14 29.35 51.74 97.97 ± 0.29 21.32 27.76 96.34 ± 0.11 28.45 48.45 97.03 ± 0.15 14.33 30.43
RD 92.01 ± 0.11 31.04 56.85 97.52 ± 0.26 19.93 24.87 97.21 ± 0.07 26.09 45.24 97.43 ± 0.07 9.98 28.87
BT 81.45 ± 0.13 39.65 54.24 95.38 ± 0.11 33.62 43.67 96.45 ± 0.14 29.47 45.38 97.65 ± 0.14 10.43 25.34
MixCode 94.32 ± 0.16 29.82 42.56 97.32 ± 0.13 12.49 18.32 96.91 ± 0.11 28.73 48.87 - - -
CodeT5 GenCode 98.85 ± 0.11 23.57 44.76 99.64 ± 0.06 12.13 19.56 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
5.3 Importance-Guided Data Selection
After constructing the training candidate (as introduced in Section 5.2), the next step is to choose the most valuable code data that can benefit the code model training. Based on our Finding drawn from the preliminary study in Section 4, GenCode chooses the simplest way to measure the importance of code data to the model – the loss value. That means, GenCode takes all the transformed code as input to the code model and records all the loss values first, then ranks all the code in descending order based on the loss values and selects the first $K$ code as the training data. Following the common setting in the data augmentation problem Dong et al. (2023a), $K$ is set as the number of original training data. In GenCode, we choose loss value as the indicator since it can be generalized to any code model and task.
6 Evaluation
We evaluate both the accuracy (Section 6.2) and robustness (Section 6.3) of trained code models. Besides, we check how does importance score affects the effectiveness of GenCode (Section 6.4).
6.1 Datasets and Models
Datasets and Models We evaluate GenCode on four code understanding tasks, bug detection Hu et al. (2019), authorship attribution Yang et al. (2022), code clone detection Yang et al. (2022) (Note that the SOTA method, MixCode, is not suitable for this task since its input is code pairs.), and problem classification Puri et al. (2021). For each code task, we fine-tune three types of models based on pre-trained code models, CodeBERT, GraphCodebERT, and CodeT5. For the model training, we follow existing works Hu et al. (2019); Wang et al. (2021); Puri et al. (2021); Yang et al. (2022); Dong et al. (2023a), to set all configurations. Specifically, we set the training epoch as 50 for all tasks and models. Regarding the optimizer, we utilize Adam Kingma and Ba (2014) with a learning rate of $10^{-3}$ for all the models. Besides, we use early stopping with a patience of 20 to mitigate overfitting. To reduce the influence of randomness, we train each model five times and report the average results along with the standard deviation. For MixCode, we follow the recommendation of the original paper and set the Mixup ratio as $\alpha=0.1$. For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  Yang et al. (2022) in our work. More details are explained in Appendix A and C.
Evaluation Metrics. In the evaluation, we assess the performance of trained models from two perspectives, namely, clean accuracy and model robustness. For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT Yang et al. (2022) and MHM Zhang et al. (2020) on code models. We do not consider the natural robustness since the existing natural robustness datasets can not fully support our studied tasks, i.e., no natural robustness datasets support authorship attribution, bug detection, and code clone detection.
Baseline Methods. We consider model training without data augmentation (No Aug), Refactor, a family of text-oriented data augmentation methods, and MixCode as baseline methods. Refactor is the widely used general data augmentation framework that has been specifically designed for code data, while MixCode is the more recent code data augmentation method with high performance in enhancing code models Dong et al. (2023a). These baselines allow us to determine if there are real improvements in the accuracy and robustness of the code model when using GenCode.
6.2 Accuracy Evaluation
Overall accuracy. Column Accuracy in Table 3 summarizes the final test accuracy of code models trained by using different code data augmentation methods including GenCode. Surprisingly, from the results, we can see that most of the existing code augmentation methods except RS and MixCode negatively affect the performance of code models. This finding is similar to the conclusion from recent code augmentation studies Bielik and Vechev (2020); Yu et al. (2022); Dong et al. (2023b) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
Then, compared to the existing data augmentation methods, the results demonstrate that GenCode consistently enhances the training and achieves the best results across all the considered datasets and models. Concretely, on average, GenCode generates models with 4.52%, 3.05%, 2.56%, and 2.11% higher accuracy than No Aug in GCJ, Refactory, Java250, and BigCloneBench, respectively. Besides, compared to the SOTA code augmentation method MixCodde, GenCode can achieve 4.1%, 2.34%, and 2.33% better results in GCJ, Refactory, and Java250 datasets, respectively. In conclusion, GenCode is effective in training accurate code understanding models.
(a) Authorship attribution
(b) Bug detection
(c) Problem classification
(d) Clone detection
Figure 2: Convergence speed of CodeBERT using different code data augmentation methods in each task.
Case study 1: Does the training epoch affect the accuracy of GenCode? Figure 2 depicts the average test accuracy at each training epoch of CodeBERT using different data augmentation methods. The results demonstrated that GenCode consistently outperforms other baseline methods during the whole training process. We can conclude that the choice of the number of training epochs will not affect the advantages of our method over other methods.
(a) Refactor
(b) SR
(c) RI
(d) RS
(e) RD
(f) BT
(g) MixCode
(h) GenCode
Figure 3: Visualization of code embeddings after dimension reduction using Principal Component Analysis (PCA). Model: CodeBERT, dataset: Refactory, task: Bug detection.
Table 4: A case study of mean values of maximum probabilities. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
Model DA method GCJ Refactory Java250 BigCloneBench
No Aug 0.0381945 0.9368535 0.9554731 0.9651272
Refactor 0.0391864 0.9643155 0.9620814 0.9654681
SR 0.0380562 0.9527686 0.9341456 0.9583812
RI 0.0371215 0.9189242 0.9327171 0.9704512
RS 0.0392987 0.9561591 0.9581767 0.9410937
RD 0.0392987 0.9723022 0.9634647 0.9631375
BT 0.0379581 0.9566747 0.9417861 0.9667154
MixCode 0.0390714 0.9687351 0.9685594 -
CodeBERT GenCode 0.0413779 0.9808122 0.9714346 0.9786421
Table 5: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data and ASR $\downarrow$ (%) on test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
GCJ Refactory Java250 BigCloneBench
Model DA method Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$
GenCode (Random) 94.74 ± 0.21 25.21 50.32 98.82 ± 0.15 27.65 36.65 98.24 ± 0.11 32.52 44.87 98.27 ± 0.19 8.01 18.12
GenCode (Min) 93.34 ± 0.32 35.66 66.67 96.13 ± 0.21 65.51 69.76 96.43 ± 0.17 46.76 53.86 97.23 ± 0.14 20.65 41.32
CodeBERT GenCode 96.23 ± 0.18 22.33 42.62 99.32 ± 0.24 22.75 29.25 98.89 ± 0.09 29.89 40.65 98.93 ± 0.15 7.12 15.83
GenCode (Random) 95.26 ± 0.11 29.97 50.63 98.93 ± 0.18 17.56 24.29 98.68 ± 0.22 20.65 38.86 98.67 ± 0.13 2.04 5.97
GenCode (Min) 93.53 ± 0.17 40.55 68.23 96.31 ± 0.24 38.73 46.98 96.72 ± 0.12 26.87 46.67 97.22 ± 0.21 11.45 21.23
GraphCodeBERT GenCode 97.93 ± 0.14 26.43 39.87 99.45 ± 0.17 15.78 20.92 99.16 ± 0.13 18.76 33.81 99.23 ± 0.09 1.86 4.62
GenCode (Random) 96.24 ± 0.21 25.03 47.61 98.99 ± 0.21 18.46 23.67 98.93 ± 0.16 25.11 42.34 98.83 ± 0.08 8.67 17.42
GenCode (Min) 94.11 ± 0.31 38.65 59.89 98.29 ± 0.12 35.67 48.59 97.03 ± 0.23 32.61 49.27 97.01 ± 0.15 18.36 25.54
CodeT5 GenCode 98.85 ± 0.16 23.57 38.76 99.64 ± 0.37 12.13 18.81 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
Case study 2: What factors contribute to the improved accuracy observed in code models when utilizing GenCode? Furthermore, we explore why models trained by GenCode have higher accuracy than models trained by using other code data augmentation models. Specifically, 1) first, we extract the code embeddings produced by each trained model using all the test data as code features. 2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) Maćkiewicz and Ratajczak (1993) algorithm. 3) After that, we visualize the reduced 2-dimensional vectors. Normally, the features extracted by a well-trained model are more clearly distributed based on the data labels. Figure 3 depicts the visualized features extracted from CodeBERT using the Refactory dataset. We can see that the code embeddings trained by using GenCode are more clearly distributed than embedding extracted by other models, i.e., the buggy code and correct code are greatly different in the figure. These results indicate that GenCode can train good code embeddings and, therefore, produce better accurate code models for code understanding tasks.
6.3 Robustness Evaluation
Overall robustness. Column MHM and ALERT in Table 3 summarize the attack success rate of code models. First, we can see that similar to the finding in Section 6.2, not all data augmentation methods can improve the robustness of code models. Specifically, in the baseline methods, only RS and MixCode have a positive impact on the robustness of code models. Considering our method, we can see that GenCode consistently improves the robustness of code models and achieves the best results in most cases. On average, compared to No Aug, models trained by GenCode have an 8.42% lower ASR (higher robustness). This result breaks the previous finding Bielik and Vechev (2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness. On the other hand, compared to the SOTA data augmentation method MixCode, GenCode can train code models with 4.90% higher robustness on average. In conclusion, GenCode can produce more robust code understanding models than other methods.
Case study 3: What factors contribute to the improved robustness observed in code models when utilizing GenCode? To explore why GenCode produces a more robust code model, we check the confidence of trained models using different code data augmentation methods on the original test data. Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data Wu et al. (2018). To do so, for each dataset and code model, we collect the maximum probability of the corrected classified test data produced by the code model first. Then, we calculate the mean values of these maximum probability scores. The results of the mean values of maximum probabilities are shown in Table 4. We can see that code models trained by GenCode are more confident in the test data than models trained using other code data augmentation methods.
6.4 Ablation Study
Even though we have preliminarily shown that selecting the code samples with higher loss values is a valid way to train good models in Section 4. It is still necessary to study to what extent the advantage can be brought from this selection strategy. Therefore, in this part, we conduct an ablation study to compare GenCode with its two variants, GenCode with data selection using minimum loss values and GenCode with random selection, GenCode (Min) and GenCode (Random), respectively.
Table 5 summarizes the accuracy and robustness of trained code models using different selection strategies. We can see GenCode consistently outperforms the two variants in both testing criteria. Selecting code samples with minimum loss values is the worst way for GenCode. Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on all results. This indicates simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods is already a good code data augmentation framework.
7 Conclusion
We proposed the first generation-and-selection-based framework, GenCode, for boosting code understanding models. The basic assumption behind GenCode is that the less confident data contributes more to the model training. Based on this assumption, GenCode utilizes atomic data augmentation operators to generate candidate training data first and selects data with maximum loss values to train the code model. Our evaluation demonstrated that GenCode can produce more accurate and robust models than MixCode with 2.92% accuracy and 4.90% robustness improvement, respectively. Besides, we analyzed how the importance score affects the effectiveness of GenCode.
In the future, we plan to 1) explore the combination of GenCode with other confidence scores, such as uncertainty metrics and active learning methods; 2) study the effectiveness of GenCode on fine-tuning open-source large language models, e.g., Llama.
Appendix A Datasets and Models
We evaluate GenCode on four code understanding tasks and three pre-trained code models as listed in Table 6. For all datasets, we follow the settings provided by the official projects for splitting the data into training, validation, and test sets.
  • Bug detection in source code understanding involves determining whether a piece of code contains bugs. This task can be thought of as a binary classification problem, where the target is to classify code as either bug-free or containing bugs. In our study, we utilize the Refactory dataset Hu et al. (2019), which is specifically designed for bug repair tasks. Refactory comprises 2,242 correct and 1,783 buggy Python programs. These programs were written by real-world undergraduate students to finish five different programming assignments.
  • Authorship attribution is the identification of the writer of a specific code fragment by inferring the unique characteristics of programmers from their published source code. This task serves important purposes, such as properly crediting a programmer for their contributions and aiding in the detection of code plagiarism. In our experiment, we employ a dataset sourced from Google Code Jam (GCJ) that is provided by the work Yang et al. (2022) for the authorship attribution task.
  • Clone detection is the task that focuses on determining whether two pieces of code are semantically identical. This task is instrumental in preventing the propagation of bugs and simplifying the process of software maintenance. To conduct our experiment, we adopt the widely acknowledged clone detection benchmark dataset BigCloneBench Svajlenko et al. (2014). This dataset offers an extensive and diverse set of code examples that are written by Java. Note that MixCode is not suitable for this task since its input is code pairs.
  • Problem classification is a typical task in source code understanding. It involves classifying the target functions of source code. Specifically, in a series of problems including detailed descriptions as well as their corresponding candidate source code, the trained model can correctly associate the source code with the specific problem that the code is trying to address. In our study, we leverage the recently released dataset Java250 Puri et al. (2021). Java250 is curated explicitly for Java program classification, containing a diverse set of 250 classification problems. Each problem within this dataset is paired with 300 different Java programs, each for solving this problem.
For the models, we study the powerful pre-trained code models.
  • CodeBERT Feng et al. (2020) is a bimodal model that is trained by leveraging data from various programming languages like C, C++, and Java, along with natural languages. Inspired by the principles of BERT Devlin et al. (2018), CodeBERT treats programs as sequences during pre-training.
  • GraphCodeBERT Guo et al. (2020) enhances the training data with data-flow information to incorporate the semantic-level structure of programs. This addition enriches the training process, enabling the model to generate more precise and context-aware code representations.
  • CodeT5 Wang et al. (2021) is a pre-trained encoder-decoder Transformer model that is designed to effectively capture the code semantics conveyed through developer-assigned identifiers.
Table 6: Details of tasks, datasets, and DNN models. #Training and #Test are the number of data for training and testing, respectively.
Dataset Language Task #Training #Test Model
Google Code Jam (GCJ) Python Authorship attribution 528 132 CodeBERT GraphCodeBERT CodeT5
Refactory Python Bug detection 3,380 423
Java250 Java Problem classification 48,000 15,000 CodeBERT GraphCodeBERT CodeT5
BigCloneBench Java Clone detection 90,102 4,000
Appendix B Atomic data augmentation operators
Code Refactoring (Refactor). Code refactoring provides a range of methods aimed at altering the syntactic structure of source code while preserving its underlying semantic information. Currently, code refactoring is being increasingly employed for code data augmentation since it can be utilized to automatically create additional source code samples. Therefore, we first chose code refactoring as one of the code data augmentation methods to generate the training candidate. Furthermore, we release all 18 atomic code refactoring operations adopted by GenCode.
Synonym Replacement (SR). The original SR is applied to text data. It augments a text by randomly selecting a certain number of words and then replacing the selected words with one of their randomly chosen synonyms. Here, synonyms are machine-generated. In contrast to text data, a program is composed of a series of statements. To augment the code data, we begin by randomly selecting a set number of statements from a program. Subsequently, each word within these selected statements is randomly replaced with one of its synonyms. All operations that may alter the syntactic structure follow the rule that preserves the original relationship between the features and the label in the code.
Random Insertion (RI). For text data, RI first randomly chooses a word in a sentence and then generates a number of synonyms of the chosen word. Lastly, RI randomly selects one of the generated synonyms and inserts it into a random position of this sentence. Generally, we repeat this operation $n$ times. In code data, we first randomly choose one of the statements and generate a number of synonyms for each word in this statement. Then, we randomly choose one word and insert one of its synonyms in a random position of the chosen statement.
Random Swap (RS). RS generally randomly chooses two different words in a sentence or two different sentences in a document and then swaps their positions. This operation is usually repeated $n$ times. For source code, we first randomly select two different statements in a program and then swap their positions, also we repeat it $n$ times.
Random Deletion (RD). RD randomly selects a certain number of words in a sentence or a certain number of sentences in a document and then deletes them. For source code, we randomly delete a certain number of words in a randomly chosen statement. Similarly, this method can preserve the original relationship between the features and the label in code data.
Back-translation (BT). BT translates the original text into another language and then translates it back to the original one to generate additional text data. For code data, similarly, we choose the English-French translation model Xie et al. (2020) bidirectionally for each statement in a program.
Not that: Different from Refactor and SR, atomic data augmentation operators including RI, RS, RD, and BT can break syntax and alter the label of the code. Despite this, we found that these code augmentation operators remain beneficial in GenCode for boosting the training of code models.
Appendix C Robustness Testing of Code Model
Robustness is an important property of DL models which indicates how the DL model can handle data with different distribution of the training data. Normally, there are two types of robustness of DL models, natural robustness and adversarial robustness. Natural robustness indicates how models perform when facing the natural distribution shift, for example, data in different environments and newly collected data. Adversarial robustness refers to the ability of models to defend the adversarial attacks. Robustness testing of code models has been addressed recently. Researchers proposed distribution shifted code (e.g., codes have the same functionality but are written by different programmers) data to study the natural robustness of code models. Besides, multiple adversarial attack methods have been proposed to test the adversarial robustness of code models, for example, Yang et al. (2022) proposed to change the variable name of code to force the model to make wrong predictions.
In this work, following existing works Yang et al. (2022), we employ adversarial attack methods proposed for code to evaluate the adversarial robustness of our trained model, namely, ALERT Yang et al. (2022) and MHM Zhang et al. (2020).
$ASR=\frac{\#Successfully\,generated\,adversarial\,data}{\#Corrected\,classified\,test\,data}\%$ (1)
In the evaluation of robustness, ASR (in Eq.( 1)) is the metric for assessing robustness. It calculates the percentage of successfully generated adversarial examples. Adversarial attacks are performed exclusively on correctly predicted test samples. A higher ASR suggests that an attack method is more effective, indicating lower robustness of the victim model.

### v2

@@
-Importance Guided Data Augmentation for Neural-Based Code Understanding
+GenCode: A Generic Data Augmentation Framework for Boosting Deep Learning-Based Code Understanding
 Abstract
-Pre-trained code models lead the era of code intelligence.
-Many models have been designed with impressive performance recently.
-However, one important problem, data augmentation for code data that automatically helps developers prepare training data lacks study in the field of code learning.
-In this paper, we introduce a general data augmentation framework, GenCode, to enhance the training of code understanding models.
-GenCode follows a generation-and-selection paradigm to prepare useful training codes.
-Specifically, it uses code transformation techniques to generate new code candidates first and then selects important ones as the training data by importance metrics.
-To evaluate the effectiveness of GenCode with a general importance metric – loss value, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5).
+Pre-trained code models lead the era of code intelligence with multiple models have been designed with impressive performance.
+However, one important problem, data augmentation for code data that automatically helps developers prepare training data lacks study in this field.
+In this paper, we introduce a generic data augmentation framework, GenCode, to enhance the training of code understanding models.
+Simply speaking, GenCode follows a generation-and-selection paradigm to prepare useful training code data.
+Specifically, it employs code transformation techniques to generate new code candidates first and then selects important ones as the training data by importance metrics.
+To evaluate the effectiveness of GenCode, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5).
 Compared to the state-of-the-art (SOTA) code augmentation method, MixCode, GenCode produces code models with 2.92% higher accuracy and 4.90% robustness on average.
-1 Introduction
+Introduction
 Deep learning (DL) for code has been one of the successful DL applications over the years.
-The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection Zhou et al.
-(2019), is commonly known as code learning.
-From the model level, many DL models specifically designed for code learning have been proposed, such as CodeBERT Feng et al.
-(2020), GraphCodeBERT Guo et al.
-(2020), and CodeT5 Wang et al.
-(2021).
-From the business perspective, multiple related products such as Copilot  [^1] https://github.com/features/copilot and Tabnine  [^2] https://www.tabnine.com have been widely used to help our daily development.
-As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  Lu et al.
-(2021) and program classification Puri et al.
-(2021), DL already achieved significantly better results than traditional SE techniques.
-However, most of the existing research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models Sun et al.
-(2022).
+The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection (Zhou et al.
+2019), is commonly known as code learning.
+From the model level, different DL models specifically designed for code learning have been proposed, such as CodeBERT (Feng et al.
+2020), GraphCodeBERT (Guo et al.
+2020), and CodeT5 (Wang et al.
+2021).
+From the commercial perspective, multiple related products such as Copilot  [^1] https://github.com/features/copilot and Tabnine  [^2] https://www.tabnine.com have been widely used to help our daily development.
+As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  (Lu et al.
+2021) and program classification (Puri et al.
+2021), DL already achieved significantly better results than traditional SE techniques.
+However, most of the existing research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models (Sun et al.
+2022).
 The challenge behind this is preparing training data for code models is heavy work.
 Although today developers can easily collect raw code data from the open-source community such as GitHub  [^3] https://github.com, data labeling is still expensive, especially for code-related tasks.
-It needs human effort with professional domain knowledge.
-Techniques can be directly used for code augmentation, such as code refactoring methods and text transformation methods.
-Unfortunately, existing studies Bielik and Vechev (2020); Yu et al.
-(2022) stated that code refactoring methods have limited advantages in boosting the training of code models.
-Besides,  Dong et al.
-(2023b) revealed that almost all the data augmentation methods cannot produce code models with good robustness.
-Moreover, a recent work, MixCode Dong et al.
-(2023a), that follows the spirit of Mixup Zhang et al.
-(2018) to do code augmentation achieved promising results, but it can be only used for classification tasks.
-Therefore, there are no effective and reliable code augmentation methods that can be generalized to different tasks.
-To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision Cubuk et al.
-(2020); Yang et al.
-(2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding.
+It needs huge human effort with professional domain knowledge.
+To tackle this problem, data augmentation is a famous technique to increase the diversity and quality of training data based only on existing labeled data.
+In the code learning field techniques such as code refactoring methods and text-oriented transformation methods can be directly used for code augmentation.
+Unfortunately, existing studies (Bielik and Vechev 2020; Yu, Wang, and Wang 2022) stated that code refactoring methods have limited advantages in the accuracy improvement of code models.
+Besides,  (Dong et al.
+2023b) revealed that code models trained using existing data augmentation methods still suffer from the low robustness issue.
+Moreover, a recent work, MixCode (Dong et al.
+2023a), that follows the spirit of Mixup (Zhang et al.
+2018) to do code augmentation achieved promising results, but it can be only used for classification tasks.
+Therefore, there are currently no effective and reliable code augmentation methods that can be generalized across different tasks.
+To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision (Cubuk et al.
+2020; Yang et al.
+2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding.
 Specifically, at each training epoch, GenCode first generates training candidates by using different code transformation methods.
-Here, we consider different types of code transformation methods such as semantic-preserving code refactoring methods and syntax-breaking text transformation methods.
+Here, we consider different types of methods such as semantic-preserving code refactoring methods and syntax-breaking text transformation methods.
 After that, GenCode selects valuable training data from the candidates with the guidance of importance metrics.
@@
 To summarize, the main contributions of this paper are:
-  • We propose the first generation-and-selection-based framework, GenCode, which automatically enhances the training data and identifies the most important data from the augmented set for boosting the code understanding.
+  • We propose the first generation-and-selection-based framework, GenCode, which automatically enriches the training data for boosting the code understanding models.
   • We empirically demonstrate that GenCode can produce more accurate and robust code understanding models compared to existing data augmentation methods and conduct an ablation study to explore how different importance scores affect the effectiveness of GenCode.
-2 Background and Related Work
-2.1 Deep Learning for Code
-Since source code can be seen as natural language or structural graph Allamanis et al.
-(2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated bug detection Dinella et al.
-(2020), automated program classification Puri et al.
-(2021), and code clone detection Wang et al.
-(2020).
-The key problem in deep learning for code is how to represent raw code to machine-readable inputs.
-Researchers designed different methods for code representation that can be roughly divided into two groups, sequence-based representation and graph-based representation.
-Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens.
-For example, given a code snippet, “def func(x, y)”, sequence-based representation transforms it to a token list as “[’def’, ’func’, ’(’, ’x’, ’y’, ’)’]”.
-After that, each token in the token list will be mapped to a pre-defined index value, e.g., [1, 25, 34, 5100, 640, 33].
-Finally, these index values will be fed into the code-embedding models for the learning.
+Background and Related Work
+Code Representation
+Code representation is a crucial technique that converts source code into a format interpretable by Deep Neural Networks (DNNs), facilitating effective learning of source code features.
+Since source code can be seen as sequences of tokens or structural graph (Allamanis et al.
+2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated program classification (Puri et al.
+2021) and code clone detection (Wang et al.
+2020).
+Existing methods for code representation can be roughly divided into two groups, sequence-based representation and graph-based representation.
+Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens (Alon et al.
+2019).
 Graph-based representation learns the structural information of code.
-Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information Dinella et al.
-(2020); Yasunaga and Liang (2020).
-For example, GraphCodeBERT combines the sequence-based representation and data flow information of code as a whole to train code models with better performance.
-In this work, we focus on both sequence-based and graph-based representation and study the powerful pre-trained code models CodeBERT, GraphCodeBERT, and CodeT5 that achieved impressive results on code understanding tasks.
-2.2 Data Augmentation for Code
-Data augmentation is a well-studied technique in the machine learning field to increase the size and diversity of training data by only modifying the existing labeled data, thus, without further human effort.
-For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym Feng et al.
-(2021).
-Recently, data augmentation gained attention from software engineering researchers.
-Various data augmentation methods designed for source code have been proposed Gao et al.
-(2020); Yefet et al.
-(2020); Allamanis et al.
-(2021); Pour et al.
-(2021); Bui et al.
-(2021); Wang et al.
-(2022); Park et al.
-(2023); Yang et al.
-(2023a), with the most general framework involving the use of code refactoring, e.g., local variable renaming, duplication, and dead store, to create more code data without breaking the semantics of the original codes.
-Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information Lacerda et al.
-(2020).
-Existing works Bielik and Vechev (2020); Yu et al.
-(2022); de Paula Rodrigues et al.
-(2023); Dong et al.
-(2023b, 2024) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models.
-For instance, a recent empirical study Dong et al.
-(2024) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation.
-Besides, the SOTA code data augmentation method, MixCode which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring.
+Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information (Dinella et al.
+2020; Yasunaga and Liang 2020).
+Unlike code representation, our work focuses on code data preprocessing and studies the powerful pre-trained code models CodeBERT, GraphCodeBERT, and CodeT5 that achieved impressive results on code understanding tasks.
+Data Augmentation for Code Understanding
+Data augmentation is a well-studied technique in the machine learning field that increases the size and diversity of training data by only modifying the existing labeled data, thus, without further human effort.
+For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym (Feng et al.
+2021).
+Recently, data augmentation has gained attention from software engineering researchers.
+Various data augmentation methods designed for source code have been proposed (Gao et al.
+2020; Yefet, Alon, and Yahav 2020; Allamanis, Jackson-Flux, and Brockschmidt 2021; Pour et al.
+2021; Bui, Yu, and Jiang 2021; Wang et al.
+2022; Park, Kim, and Han 2023; Yang et al.
+2023a), with the most general framework involving the use of code refactoring, e.g., local variable renaming and dead store, to create more code data without breaking the semantics of the original codes.
+Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information (Lacerda et al.
+2020).
+Existing works (Bielik and Vechev 2020; Yu, Wang, and Wang 2022; de Paula Rodrigues, Braga, and Dahab 2023; Dong et al.
+2023b, 2024) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models.
+For instance, a recent empirical study (Dong et al.
+2024) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation.
+Besides, the SOTA code data augmentation method, MixCode, which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring.
 However, as a variant of Mixup specifically designed for code data, MixCode is limited to applications in code classification tasks, as it necessitates the label with one-hot encoding.
 Compared to existing works, we propose a simple yet effective code augmentation framework that is not only unaffected by the dataset and model constraints but also applicable to both classification and non-classification tasks.
-3 Definition
+Definition
 We first define the problem of data augmentation for code models.
-Definition 3.1 (Program Code [^4] For simplicity, we use code in the paper. ).
-A code is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: $content$ contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
-Definition 3.2 (Code Model).
+Definition 1.
+A program code  [^4] For simplicity, we use code in the paper.
+is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: $content$ contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
+Definition 2.
 A code model is defined as a 3-tuple $M=\left(parameter,\mathcal{P},Y\right)$, where $parameter$ contains all the weights in code model $M$, $\mathcal{P}$ is the input code, and $Y$ is the label of $\mathcal{P}$ given a specific downstream code task.
-Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 Wang et al. (2021).
-Definition 3.3 (Code Augmentation).
-Given a code $P$, a code transformation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)$.
+Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 (Wang et al. 2021).
+Definition 3.
+Given a code $P$, a code transformation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)\&\mathcal{P^{\prime}}.content\neq\mathcal{P}.content$.
 Here, code transformation methods can be any method that can change the content of program statements such as text data transformation methods (e.g., synonym replacement and code refactoring methods (e.g., dead if statement adding).
-Definition 3.4 (Semantic-preserving/breaking Code Data Augmentation).
+Definition 4.
 Given a code model $M$, a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, semantic-preserving (breaking) code augmentation indicates $\mathcal{P}.se=\mathcal{P^{\prime}}.se$ $M(\mathcal{P})=M(\mathcal{P^{\prime}})$ ($\mathcal{P}.se\neq\mathcal{P^{\prime}}.se$ $M(\mathcal{P})\neq M(\mathcal{P^{\prime}})$).
-Definition 3.5 (Syntax-preserving/breaking Code Augmentation).
+Definition 5.
 Given a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, syntax-preserving (breaking) code augmentation indicates $\mathcal{P}.sy=\mathcal{P^{\prime}}.sy$ ($\mathcal{P}.sy\neq\mathcal{P^{\prime}}.sy$).
-4 Assumption and Preliminary Study
+Assumption and Preliminary Study
 GenCode selects the most valuable data for the code model training.
@@
 To check whether this assumption is valid or not, we conducted a preliminary study to explore if there is a correlation between the accuracy of the trained code model and the confidence of the model on the training data.
-Specifically, the study is conducted on the Google Code Jam (GCJ) Yang et al.
-(2022) dataset and CodeBERT pre-trained code model.
+Specifically, the study is conducted on the Google Code Jam (GCJ) (Yang et al.
+2022) dataset and CodeBERT pre-trained code model.
 The detailed steps are as follows:
-Step 1. We initialize the code model and use code augmentation methods to generate additional data as the training candidate set. The code augmentation methods will be introduced in Section 5.2.
+Step 1. We initialize the code model and use code augmentation methods to generate additional data as the training candidate set. The code augmentation methods will be introduced in a later section.
 Step 2.
@@
 We repeat this selection 100 times and collect 100 groups of training data.
-After that, for each group of training data, we compute and record their average loss values.
+After that, we compute and record the average loss values for each group of training data.
 Step 3.
@@
 We then test each code model using the original test data and save 100 accuracies.
-Step 4. We use the Pearson Correlation Coefficient (PCC) Schober et al. (2018) metric to check the correlation between the loss values and the accuracies.
-Table 1: Correlation between loss values and code model accuracy.
+Step 4. We use the Pearson Correlation Coefficient (PCC) (Schober, Boer, and Schwarte 2018) metric to check the correlation between the loss values and the accuracies.
+Table 1: Correlation between loss values and code model accuracy. The P-value less than 0.01 is highlighted with a blue background.
 Coefficient P-value
@@
 From the initialized model to the well-trained model, this correlation becomes weaker but still exists.
-Therefore, we can propose an effective and reliable data augmentation framework to boost the training of code models based on this phenomenon.
+We conjecture that when the model is in the early stage, most of the data that have high loss values are useful for the model training.
+However, when training proceeds, the data that still have high loss values are the model itself cannot learn.
+Thus, there will be a correlation drop from the early epoch to the later ones.
+Based on this phenomenon, we can propose an effective and reliable data augmentation framework to boost the training of code models based on this phenomenon.
 Finding: There is a positive correlation between the loss values of code data and the accuracy of code models trained by using these data.
 When the code model is randomly initialized, this correlation is significant (with a P-value less than 0.01).
-5 Methodology
-5.1 Overview
+Methodology
+Overview
 The goal of GenCode is to produce code models with better performance on test data by training models on carefully generated data.
@@
 Algorithm 1 shows how GenCode works.
-In each training epoch, GenCode automatically generates more code data (Line 7 to Line 9, Definition 3.3) and then selects the most important ones (Line 10 to Line 13) to train the code model (Line 14).
+In each training epoch, GenCode automatically generates more code data (Line 7 to Line 9, Definition 3) and then selects the most important ones (Line 10 to Line 13) to train the code model (Line 14).
 Here, the data importance is measured by the loss values produced by the code model which is the most common indicator to check how the model performs on the data.
@@
 15:return $M$
-5.2 Search Space
-Even though the code has strict syntax constraints, recent work Dong et al.
-(2023b) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training.
-To enhance the diversity of source code within the feature space, therefore, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 3.4 and Definition 3.5).
+Search Space
+Even though the code has strict syntax constraints, recent work (Dong et al.
+2023b) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training.
+To enhance the diversity of source code within the feature space, therefore, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 4 and Definition 5).
 GenCode adapts code refactoring methods as semantic/syntax-preserving methods.
@@
 As depicted in this example, the third and fourth lines have been interchanged.
-Appendix B detailed introduces the specifics of these atomic data augmentation operators.
 Table 2: Atomic data augmentation operators adopted in GenCode.
@@
 Filed enhancement
-Synonym replacement Random insertion Random swap Random deletion Back-translation
+Synonym replacement (SR) Random insertion (RI) Random swap (RS Random deletion (RD) Back-translation (BT)
 Arguments adding
@@
 GCJ Refactory Java250 BigCloneBench
-Model DA method Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$
+Model DA method Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$
 No Aug 90.96 ± 0.07 33.31 56.13 96.21 ± 0.13 36.31 42.31 96.36 ± 0.09 40.55 53.37 96.91 ± 0.12 9.74 24.61
@@
 CodeT5 GenCode 98.85 ± 0.11 23.57 44.76 99.64 ± 0.06 12.13 19.56 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
-5.3 Importance-Guided Data Selection
-After constructing the training candidate (as introduced in Section 5.2), the next step is to choose the most valuable code data that can benefit the code model training.
-Based on our Finding drawn from the preliminary study in Section 4, GenCode chooses the simplest way to measure the importance of code data to the model – the loss value.
+Importance-Guided Data Selection
+After constructing the training candidate, we then choose the most valuable code data that can benefit the code model training.
+Based on our Finding drawn from the preliminary study, GenCode chooses the simplest way to measure the importance of code data to the model – the loss value.
 That means, GenCode takes all the transformed code as input to the code model and records all the loss values first, then ranks all the code in descending order based on the loss values and selects the first $K$ code as the training data.
-Following the common setting in the data augmentation problem Dong et al.
-(2023a), $K$ is set as the number of original training data.
-In GenCode, we choose loss value as the indicator since it can be generalized to any code model and task.
-6 Evaluation
-We evaluate both the accuracy (Section 6.2) and robustness (Section 6.3) of trained code models. Besides, we check how does importance score affects the effectiveness of GenCode (Section 6.4).
-6.1 Datasets and Models
-Datasets and Models We evaluate GenCode on four code understanding tasks, bug detection Hu et al.
-(2019), authorship attribution Yang et al.
-(2022), code clone detection Yang et al.
-(2022) (Note that the SOTA method, MixCode, is not suitable for this task since its input is code pairs.), and problem classification Puri et al.
-(2021).
+Following the common setting in the data augmentation problem (Dong et al.
+2023a), $K$ is set as the number of original training data.
+GenCode chooses loss value as the indicator since it can be generalized to any code model and task.
+Evaluation
+We evaluate both the accuracy and robustness of trained code models. Besides, we check how does importance score affects the effectiveness of GenCode.
+Datasets and Models
+Datasets and Models We evaluate GenCode on four code understanding tasks, bug detection (Hu et al.
+2019), authorship attribution (Yang et al.
+2022), code clone detection (Yang et al.
+2022) (Note that, MixCode is not suitable for this task since its input is code pairs.), and problem classification (Puri et al.
+2021).
 For each code task, we fine-tune three types of models based on pre-trained code models, CodeBERT, GraphCodebERT, and CodeT5.
-For the model training, we follow existing works Hu et al.
-(2019); Wang et al.
-(2021); Puri et al.
-(2021); Yang et al.
-(2022); Dong et al.
-(2023a), to set all configurations.
+For the model training, we follow existing works (Hu et al.
+2019; Wang et al.
+2021; Puri et al.
+2021; Yang et al.
+2022; Dong et al.
+2023a), to set all configurations.
 Specifically, we set the training epoch as 50 for all tasks and models.
-Regarding the optimizer, we utilize Adam Kingma and Ba (2014) with a learning rate of $10^{-3}$ for all the models.
+Regarding the optimizer, we utilize Adam (Kingma and Ba 2014) with a learning rate of $10^{-3}$ for all the models.
 Besides, we use early stopping with a patience of 20 to mitigate overfitting.
@@
 For MixCode, we follow the recommendation of the original paper and set the Mixup ratio as $\alpha=0.1$.
-For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  Yang et al.
-(2022) in our work.
-More details are explained in Appendix A and C.
+For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  (Yang et al.
+2022).
 Evaluation Metrics.
 In the evaluation, we assess the performance of trained models from two perspectives, namely, clean accuracy and model robustness.
-For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT Yang et al.
-(2022) and MHM Zhang et al.
-(2020) on code models.
+For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT (Yang et al.
+2022), which uses variable renaming while considering natural semantics, and MHM (Zhang et al.
+2020), which also randomly replaces variable names but does not account for natural semantics.
 We do not consider the natural robustness since the existing natural robustness datasets can not fully support our studied tasks, i.e., no natural robustness datasets support authorship attribution, bug detection, and code clone detection.
 Baseline Methods.
-We consider model training without data augmentation (No Aug), Refactor, a family of text-oriented data augmentation methods, and MixCode as baseline methods.
-Refactor is the widely used general data augmentation framework that has been specifically designed for code data, while MixCode is the more recent code data augmentation method with high performance in enhancing code models Dong et al.
-(2023a).
-These baselines allow us to determine if there are real improvements in the accuracy and robustness of the code model when using GenCode.
-6.2 Accuracy Evaluation
+We consider model training without data augmentation (No Aug), Refactor (Yu, Wang, and Wang 2022), a family of text-oriented data augmentation methods (Dong et al.
+2023b), and MixCode as baseline methods.
+Refactor, which keeps code syntax rules, is a widely used data augmentation framework specifically designed for code data.
+In contrast, text-oriented code data augmentation methods, which slightly break syntax rules, have recently proven more effective than Refactor.
+Additionally, MixCode is a more recent and high-performing code data augmentation method that significantly enhances code models.
+Accuracy Evaluation
 Overall accuracy.
@@
 Surprisingly, from the results, we can see that most of the existing code augmentation methods except RS and MixCode negatively affect the performance of code models.
-This finding is similar to the conclusion from recent code augmentation studies Bielik and Vechev (2020); Yu et al.
-(2022); Dong et al.
-(2023b) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
+This finding is similar to the conclusion from recent code augmentation studies (Bielik and Vechev 2020; Yu, Wang, and Wang 2022; Dong et al.
+2023b) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
 Then, compared to the existing data augmentation methods, the results demonstrate that GenCode consistently enhances the training and achieves the best results across all the considered datasets and models.
 Concretely, on average, GenCode generates models with 4.52%, 3.05%, 2.56%, and 2.11% higher accuracy than No Aug in GCJ, Refactory, Java250, and BigCloneBench, respectively.
-Besides, compared to the SOTA code augmentation method MixCodde, GenCode can achieve 4.1%, 2.34%, and 2.33% better results in GCJ, Refactory, and Java250 datasets, respectively.
-In conclusion, GenCode is effective in training accurate code understanding models.
+Besides, compared to MixCode, GenCode can achieve 4.1%, 2.34%, and 2.33% better results in GCJ, Refactory, and Java250 datasets, respectively.
+Statistical analysis.
+We conducted statistical tests, including the T-Test and Wilcoxon signed-rank test, across all pre-trained models.
+The results show that GenCode significantly (with p-value $<$ 0.05) outperforms other methods including MixCode.
+In conclusion, GenCode is effective in training accurate code models.
 (a) Authorship attribution
@@
 Figure 2: Convergence speed of CodeBERT using different code data augmentation methods in each task.
-Case study 1: Does the training epoch affect the accuracy of GenCode?
+Convergence speed analysis.
 Figure 2 depicts the average test accuracy at each training epoch of CodeBERT using different data augmentation methods.
@@
 Figure 3: Visualization of code embeddings after dimension reduction using Principal Component Analysis (PCA). Model: CodeBERT, dataset: Refactory, task: Bug detection.
-Table 4: A case study of mean values of maximum probabilities.
+Considering Figure 3, we can see that the embedding produced by GenCode is clearly distributed compared to other methods.
+Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Figure 3.
+The results are shown in the following Table.
+The results confirm that GenCode can better distinguish the two classes.
+Interestingly, we found that MixCode has the smallest distance result.
+That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries.
+This phenomenon is similar to the findings in (Verma et al.
+2019), where the representations obtained by interpolating hidden states have two desirable properties: the class-representations are flattened into a minimal amount of directions of variation, and all points in between these flat representations, most unobserved during training and off the data manifold, are assigned low-confidence predictions.
+Table 4: Average distance between the data points from different classes. The best is highlighted with a blue background, and the worst is highlighted with a gray background.
+DA method Refactor SR RI RS
+Average distance 6.02 8.05 7.02 8.06
+DA method RD BT MixCode GenCode
+Average distance 7.03 6.00 4.05 8.78
+Table 5: Mean values of maximum probabilities.
 The best results are highlighted in blue.
@@
 GCJ Refactory Java250 BigCloneBench
-Model DA method Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$
+Model DA method Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$
 GenCode (Random) 94.74 ± 0.21 25.21 50.32 98.82 ± 0.15 27.65 36.65 98.24 ± 0.11 32.52 44.87 98.27 ± 0.19 8.01 18.12
@@
 CodeT5 GenCode 98.85 ± 0.16 23.57 38.76 99.64 ± 0.37 12.13 18.81 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
-Case study 2: What factors contribute to the improved accuracy observed in code models when utilizing GenCode?
-Furthermore, we explore why models trained by GenCode have higher accuracy than models trained by using other code data augmentation models.
+Visualizing principal component analysis.
+Furthermore, we explore why models trained by GenCode have higher accuracy than models trained using other methods by analyzing the code representation capability of trained code embeddings.
 Specifically, 1) first, we extract the code embeddings produced by each trained model using all the test data as code features.
-2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) Maćkiewicz and Ratajczak (1993) algorithm.
+2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) (Maćkiewicz and Ratajczak 1993) algorithm.
 3) After that, we visualize the reduced 2-dimensional vectors.
@@
 These results indicate that GenCode can train good code embeddings and, therefore, produce better accurate code models for code understanding tasks.
-6.3 Robustness Evaluation
+Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Figure.
+3.
+The results from Table 4 confirm that GenCode can better distinguish the two classes.
+Interestingly, we found that MixCode has the smallest distance result.
+That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries.
+This phenomenon is similar to the findings in (Verma et al.
+2019), where interpolating hidden state representations yields two beneficial properties: 1) class representations are flattened into a minimal number of directions of variation, and 2) all points between these flat representations are assigned low-confidence predictions.
+Robustness Evaluation
 Overall robustness.
-Column MHM and ALERT in Table 3 summarize the attack success rate of code models.
-First, we can see that similar to the finding in Section 6.2, not all data augmentation methods can improve the robustness of code models.
+Column MHM and ALERT in Table 3 summarize the attack success rate of adversarial attack methods against code models.
+First, we can see that similar to the observation in Accuracy Evaluation, not all data augmentation methods can improve the robustness of code models.
+This finding is consistent with the existing works (Dong et al.
+2023b).
 Specifically, in the baseline methods, only RS and MixCode have a positive impact on the robustness of code models.
-Considering our method, we can see that GenCode consistently improves the robustness of code models and achieves the best results in most cases.
+Considering our method, GenCode consistently improves the robustness of code models and achieves the best results in 21 out of 24 cases.
 On average, compared to No Aug, models trained by GenCode have an 8.42% lower ASR (higher robustness).
-This result breaks the previous finding Bielik and Vechev (2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness.
-On the other hand, compared to the SOTA data augmentation method MixCode, GenCode can train code models with 4.90% higher robustness on average.
-In conclusion, GenCode can produce more robust code understanding models than other methods.
-Case study 3: What factors contribute to the improved robustness observed in code models when utilizing GenCode?
+This result breaks the previous finding (Bielik and Vechev 2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness.
+On the other hand, compared to MixCode, GenCode can train code models with 4.90% higher robustness on average.
+Statistical analysis.
+We also conduct statistical tests using T-Test and Wilcoxon signed-rank test.
+When comparing existing data augmentation methods, GenCode consistently exhibits a p-value $<$ 0.05 across all pre-trained models.
+In conclusion, GenCode can produce more adversarial robust code understanding models than other methods.
+Maximum probability analysis.
 To explore why GenCode produces a more robust code model, we check the confidence of trained models using different code data augmentation methods on the original test data.
-Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data Wu et al.
-(2018).
+Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data (Wu et al.
+2018).
 To do so, for each dataset and code model, we collect the maximum probability of the corrected classified test data produced by the code model first.
@@
 We can see that code models trained by GenCode are more confident in the test data than models trained using other code data augmentation methods.
-6.4 Ablation Study
-Even though we have preliminarily shown that selecting the code samples with higher loss values is a valid way to train good models in Section 4.
-It is still necessary to study to what extent the advantage can be brought from this selection strategy.
-Therefore, in this part, we conduct an ablation study to compare GenCode with its two variants, GenCode with data selection using minimum loss values and GenCode with random selection, GenCode (Min) and GenCode (Random), respectively.
+As a result, GenCode trained code models are harder to attack.
+Ablation Study
+Even though we have preliminarily shown that selecting the code samples with higher loss values is a valid way to train good models, it is still necessary to study to what extent the advantage can be brought from this selection strategy.
+Therefore, in this part, we conduct an ablation study to compare GenCode with its two variants, GenCode with data selection using minimum loss values - GenCode (Min) and random selection - GenCode (Random), respectively.
 Table 6 summarizes the accuracy and robustness of trained code models using different selection strategies.
-We can see GenCode consistently outperforms the two variants in both testing criteria.
+We can see GenCode consistently outperforms the two variants under all testing criteria.
 Selecting code samples with minimum loss values is the worst way for GenCode.
-Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on all results.
-This indicates simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods is already a good code data augmentation framework.
-7 Conclusion
-We proposed the first generation-and-selection-based framework, GenCode, for boosting code understanding models.
+Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on the results.
+This phenomenon indicates simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods is already a good code data augmentation method.
+Discussion
+Even though the accuracy of our studied code models already achieved relatively high accuracy in these tasks, i.e., most of them have more than 90% accuracy, their robustness is still low.
+Regarding the large language models (LLMs), we have checked the literature and found that the clone detection task is still challenging for LLMs, e.g., ChatGPT has only 0.784 Precision for Java-Java pair detection (Khajezade et al.
+2024).
+For other program understanding tasks that have a large number of classes, e.g., the Java250 dataset used in our experiments has 250 classes.
+It is also hard to use open-source LLMs such as ChatGPT to predict the label since we need to add many label guidance in the prompt.
+Conclusion
+We propose the first generation-and-selection-based framework, GenCode, to boost code understanding models.
 The basic assumption behind GenCode is that the less confident data contributes more to the model training.
 Based on this assumption, GenCode utilizes atomic data augmentation operators to generate candidate training data first and selects data with maximum loss values to train the code model.
-Our evaluation demonstrated that GenCode can produce more accurate and robust models than MixCode with 2.92% accuracy and 4.90% robustness improvement, respectively.
-Besides, we analyzed how the importance score affects the effectiveness of GenCode.
+Our evaluation demonstrates that GenCode can produce more accurate and robust models than MixCode with 2.92% accuracy and 4.90% robustness improvement, respectively.
+Besides, we analyze how the importance score affects the effectiveness of GenCode.
 In the future, we plan to 1) explore the combination of GenCode with other confidence scores, such as uncertainty metrics and active learning methods; 2) study the effectiveness of GenCode on fine-tuning open-source large language models, e.g., Llama.
-Appendix A Datasets and Models
-We evaluate GenCode on four code understanding tasks and three pre-trained code models as listed in Table 6.
-For all datasets, we follow the settings provided by the official projects for splitting the data into training, validation, and test sets.
-  • Bug detection in source code understanding involves determining whether a piece of code contains bugs.
-This task can be thought of as a binary classification problem, where the target is to classify code as either bug-free or containing bugs.
-In our study, we utilize the Refactory dataset Hu et al.
-(2019), which is specifically designed for bug repair tasks.
-Refactory comprises 2,242 correct and 1,783 buggy Python programs.
-These programs were written by real-world undergraduate students to finish five different programming assignments.
-  • Authorship attribution is the identification of the writer of a specific code fragment by inferring the unique characteristics of programmers from their published source code.
-This task serves important purposes, such as properly crediting a programmer for their contributions and aiding in the detection of code plagiarism.
-In our experiment, we employ a dataset sourced from Google Code Jam (GCJ) that is provided by the work Yang et al.
-(2022) for the authorship attribution task.
-  • Clone detection is the task that focuses on determining whether two pieces of code are semantically identical.
-This task is instrumental in preventing the propagation of bugs and simplifying the process of software maintenance.
-To conduct our experiment, we adopt the widely acknowledged clone detection benchmark dataset BigCloneBench Svajlenko et al.
-(2014).
-This dataset offers an extensive and diverse set of code examples that are written by Java.
-Note that MixCode is not suitable for this task since its input is code pairs.
-  • Problem classification is a typical task in source code understanding.
-It involves classifying the target functions of source code.
-Specifically, in a series of problems including detailed descriptions as well as their corresponding candidate source code, the trained model can correctly associate the source code with the specific problem that the code is trying to address.
-In our study, we leverage the recently released dataset Java250 Puri et al.
-(2021).
-Java250 is curated explicitly for Java program classification, containing a diverse set of 250 classification problems.
-Each problem within this dataset is paired with 300 different Java programs, each for solving this problem.
-For the models, we study the powerful pre-trained code models.
-  • CodeBERT Feng et al.
-(2020) is a bimodal model that is trained by leveraging data from various programming languages like C, C++, and Java, along with natural languages.
-Inspired by the principles of BERT Devlin et al.
-(2018), CodeBERT treats programs as sequences during pre-training.
-  • GraphCodeBERT Guo et al.
-(2020) enhances the training data with data-flow information to incorporate the semantic-level structure of programs.
-This addition enriches the training process, enabling the model to generate more precise and context-aware code representations.
-  • CodeT5 Wang et al. (2021) is a pre-trained encoder-decoder Transformer model that is designed to effectively capture the code semantics conveyed through developer-assigned identifiers.
-Table 6: Details of tasks, datasets, and DNN models. #Training and #Test are the number of data for training and testing, respectively.
-Dataset Language Task #Training #Test Model
-Google Code Jam (GCJ) Python Authorship attribution 528 132 CodeBERT GraphCodeBERT CodeT5
-Refactory Python Bug detection 3,380 423
-Java250 Java Problem classification 48,000 15,000 CodeBERT GraphCodeBERT CodeT5
-BigCloneBench Java Clone detection 90,102 4,000
-Appendix B Atomic data augmentation operators
-Code Refactoring (Refactor).
-Code refactoring provides a range of methods aimed at altering the syntactic structure of source code while preserving its underlying semantic information.
-Currently, code refactoring is being increasingly employed for code data augmentation since it can be utilized to automatically create additional source code samples.
-Therefore, we first chose code refactoring as one of the code data augmentation methods to generate the training candidate.
-Furthermore, we release all 18 atomic code refactoring operations adopted by GenCode.
-Synonym Replacement (SR).
-The original SR is applied to text data.
-It augments a text by randomly selecting a certain number of words and then replacing the selected words with one of their randomly chosen synonyms.
-Here, synonyms are machine-generated.
-In contrast to text data, a program is composed of a series of statements.
-To augment the code data, we begin by randomly selecting a set number of statements from a program.
-Subsequently, each word within these selected statements is randomly replaced with one of its synonyms.
-All operations that may alter the syntactic structure follow the rule that preserves the original relationship between the features and the label in the code.
-Random Insertion (RI).
-For text data, RI first randomly chooses a word in a sentence and then generates a number of synonyms of the chosen word.
-Lastly, RI randomly selects one of the generated synonyms and inserts it into a random position of this sentence.
-Generally, we repeat this operation $n$ times.
-In code data, we first randomly choose one of the statements and generate a number of synonyms for each word in this statement.
-Then, we randomly choose one word and insert one of its synonyms in a random position of the chosen statement.
-Random Swap (RS).
-RS generally randomly chooses two different words in a sentence or two different sentences in a document and then swaps their positions.
-This operation is usually repeated $n$ times.
-For source code, we first randomly select two different statements in a program and then swap their positions, also we repeat it $n$ times.
-Random Deletion (RD).
-RD randomly selects a certain number of words in a sentence or a certain number of sentences in a document and then deletes them.
-For source code, we randomly delete a certain number of words in a randomly chosen statement.
-Similarly, this method can preserve the original relationship between the features and the label in code data.
-Back-translation (BT).
-BT translates the original text into another language and then translates it back to the original one to generate additional text data.
-For code data, similarly, we choose the English-French translation model Xie et al.
-(2020) bidirectionally for each statement in a program.
-Not that: Different from Refactor and SR, atomic data augmentation operators including RI, RS, RD, and BT can break syntax and alter the label of the code.
-Despite this, we found that these code augmentation operators remain beneficial in GenCode for boosting the training of code models.
-Appendix C Robustness Testing of Code Model
-Robustness is an important property of DL models which indicates how the DL model can handle data with different distribution of the training data.
-Normally, there are two types of robustness of DL models, natural robustness and adversarial robustness.
-Natural robustness indicates how models perform when facing the natural distribution shift, for example, data in different environments and newly collected data.
-Adversarial robustness refers to the ability of models to defend the adversarial attacks.
-Robustness testing of code models has been addressed recently.
-Researchers proposed distribution shifted code (e.g., codes have the same functionality but are written by different programmers) data to study the natural robustness of code models.
-Besides, multiple adversarial attack methods have been proposed to test the adversarial robustness of code models, for example, Yang et al.
-(2022) proposed to change the variable name of code to force the model to make wrong predictions.
-In this work, following existing works Yang et al.
-(2022), we employ adversarial attack methods proposed for code to evaluate the adversarial robustness of our trained model, namely, ALERT Yang et al.
-(2022) and MHM Zhang et al.
-(2020).
-$ASR=\frac{\#Successfully\,generated\,adversarial\,data}{\#Corrected\,classified\,test\,data}\%$ (1)
-In the evaluation of robustness, ASR (in Eq.( 1)) is the metric for assessing robustness.
-It calculates the percentage of successfully generated adversarial examples.
-Adversarial attacks are performed exclusively on correctly predicted test samples.
-A higher ASR suggests that an attack method is more effective, indicating lower robustness of the victim model.

### v3

@@
 Abstract
-Pre-trained code models lead the era of code intelligence with multiple models have been designed with impressive performance.
+Pre-trained code models lead the era of code intelligence, with multiple models designed with impressive performance.
 However, one important problem, data augmentation for code data that automatically helps developers prepare training data lacks study in this field.
@@
 Simply speaking, GenCode follows a generation-and-selection paradigm to prepare useful training code data.
-Specifically, it employs code transformation techniques to generate new code candidates first and then selects important ones as the training data by importance metrics.
-To evaluate the effectiveness of GenCode, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5).
-Compared to the state-of-the-art (SOTA) code augmentation method, MixCode, GenCode produces code models with 2.92% higher accuracy and 4.90% robustness on average.
-Introduction
+Specifically, it employs code augmentation techniques to generate new code candidates first and then identifies important ones as the training data by influence scores.
+To evaluate the effectiveness of GenCode, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5) and two recent released code-specific Large Language Models (LLMs) (e.g., Qwen2.5-Coder).
+Compared to the state-of-the-art (SOTA) code augmentation method MixCode, GenCode produces pre-trained code models with 2.92% higher accuracy and 4.90% adversarial robustness on average.
+For code-specific LLMs, GenCode achieves an average improvement of 0.93% in accuracy and 0.98% in natural robustness.
+1 Introduction
 Deep learning (DL) for code has been one of the successful DL applications over the years.
-The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection (Zhou et al.
-2019), is commonly known as code learning.
-From the model level, different DL models specifically designed for code learning have been proposed, such as CodeBERT (Feng et al.
-2020), GraphCodeBERT (Guo et al.
-2020), and CodeT5 (Wang et al.
-2021).
+The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection Zhou et al.
+(2019), is commonly known as code learning.
+From the model level, different DL models specifically designed for code learning have been proposed, such as CodeBERT Feng et al.
+(2020), CodeT5 Wang et al.
+(2021), and Qwen2.5-Coder Hui et al.
+(2024).
 From the commercial perspective, multiple related products such as Copilot  [^1] https://github.com/features/copilot and Tabnine  [^2] https://www.tabnine.com have been widely used to help our daily development.
-As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  (Lu et al.
-2021) and program classification (Puri et al.
-2021), DL already achieved significantly better results than traditional SE techniques.
-However, most of the existing research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models (Sun et al.
-2022).
+As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  Lu et al.
+(2021) and program classification Puri et al.
+(2021), DL has already achieved significantly better results than traditional SE techniques.
+However, most of the existing code learning research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models Sun et al.
+(2022).
 The challenge behind this is preparing training data for code models is heavy work.
 Although today developers can easily collect raw code data from the open-source community such as GitHub  [^3] https://github.com, data labeling is still expensive, especially for code-related tasks.
-It needs huge human effort with professional domain knowledge.
-To tackle this problem, data augmentation is a famous technique to increase the diversity and quality of training data based only on existing labeled data.
-In the code learning field techniques such as code refactoring methods and text-oriented transformation methods can be directly used for code augmentation.
-Unfortunately, existing studies (Bielik and Vechev 2020; Yu, Wang, and Wang 2022) stated that code refactoring methods have limited advantages in the accuracy improvement of code models.
-Besides,  (Dong et al.
-2023b) revealed that code models trained using existing data augmentation methods still suffer from the low robustness issue.
-Moreover, a recent work, MixCode (Dong et al.
-2023a), that follows the spirit of Mixup (Zhang et al.
-2018) to do code augmentation achieved promising results, but it can be only used for classification tasks.
-Therefore, there are currently no effective and reliable code augmentation methods that can be generalized across different tasks.
-To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision (Cubuk et al.
-2020; Yang et al.
-2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding.
-Specifically, at each training epoch, GenCode first generates training candidates by using different code transformation methods.
-Here, we consider different types of methods such as semantic-preserving code refactoring methods and syntax-breaking text transformation methods.
-After that, GenCode selects valuable training data from the candidates with the guidance of importance metrics.
-The important metric used in GenCode is simply defined as the loss value of code data produced by the code model.
+It requires huge human effort and professional domain knowledge.
+To tackle this problem, data augmentation is a famous technique that increases the diversity and quality of training data based only on existing labeled data.
+In the code learning field techniques such as code refactoring methods and text-oriented data augmentation methods can be directly used for code augmentation.
+Unfortunately, existing studies Bielik and Vechev (2020); Yu et al.
+(2022) stated that code refactoring methods have limited advantages in the accuracy improvement of code models.
+Besides, the recent work Dong et al.
+(2025) revealed that code models trained using existing data augmentation methods still suffer from the low robustness issue.
+Therefore, exploring more effective and reliable code augmentation methods is an emerging task.
+To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision Cubuk et al.
+(2020); Yang et al.
+(2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding.
+Specifically, at each training epoch, GenCode first generates training candidates by using different code augmentation methods.
+Here, we consider different types of methods such as semantic-preserving code refactoring methods and syntax-breaking text augmentation methods.
+After that, GenCode selects valuable training data from the candidates with the guidance of influence scores.
+The influence score used in GenCode is simply defined as the loss value of code data produced by the code model.
 Finally, the selected samples are used for the model training.
 We conducted comprehensive experiments to study how GenCode can improve the accuracy and robustness of code models.
-The experimental results demonstrated that, compared to the SOTA method, MixCode, GenCode can produce code models with a 2.92% higher accuracy and 4.90% better robustness on average.
+Based on that, we answer the following three research questions:
+RQ1:How effective is GenCode in producing accurate code models?
+The results demonstrate that, on average, GenCode can enhance the accuracy of pre-trained code models by up to 3.63% and 2.92% compared to the training without using data augmentation and the SOTA method MixCode, respectively.
+For code-specific LLMs, Gencode achieves average improvements of 0.93% for StarCoder2 and 0.84% for Qwen2.5-Coder.
+Notably, our statistical tests show GenCode consistently outperforms baseline methods in terms of accuracy improvement.
+RQ2: How effective is GenCode in producing robust code models?
+The results demonstrate that GenCode produces more robust pre-trained code models compared to existing data augmentation methods.
+Specifically, GenCode achieves an average robustness improvement of 8.42% over models trained without data augmentation and 4.90% over models trained with MixCode.
+However, for code-specific LLMs, the robustness gains are modest, with improvements of less than 1%.
+RQ3: How does the influence score affect the effectiveness of GenCode?
+The results suggest that using the influence score based on maximum loss values is the most effective strategy among the considered variants, including minimum loss value selection and random selection.
+Besides, even GenCode with random selection outperforms existing code augmentation methods.
+RQ4: How does the data selection method affect the effectiveness of GenCode?
+Experimental results indicate that GenCode achieves superior performance in terms of accuracy improvement.
+However, the gradient-based method outperforms GenCode and other data selection methods in enhancing model robustness.
 To summarize, the main contributions of this paper are:
-  • We propose the first generation-and-selection-based framework, GenCode, which automatically enriches the training data for boosting the code understanding models.
-  • We empirically demonstrate that GenCode can produce more accurate and robust code understanding models compared to existing data augmentation methods and conduct an ablation study to explore how different importance scores affect the effectiveness of GenCode.
-Background and Related Work
-Code Representation
+  • We propose the first generation-and-selection-based framework, GenCode  [^4] https://github.com/zemingd/GenCode, which automatically enriches the training data for boosting the code understanding models.
+  • We empirically demonstrate that GenCode can produce more accurate and robust code understanding models compared to existing data augmentation methods.
+  • We conduct an ablation study to explore how different influence scores and data selection methods affect the effectiveness of GenCode.
+2 Background
+In this section, we introduce the necessary knowledge that is needed to understand our work, code representation, data augmentation for code understanding, and robustness testing of code model.
+2.1 Code Representation
 Code representation is a crucial technique that converts source code into a format interpretable by Deep Neural Networks (DNNs), facilitating effective learning of source code features.
-Since source code can be seen as sequences of tokens or structural graph (Allamanis et al.
-2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated program classification (Puri et al.
-2021) and code clone detection (Wang et al.
-2020).
-Existing methods for code representation can be roughly divided into two groups, sequence-based representation and graph-based representation.
-Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens (Alon et al.
-2019).
+Since source code can be seen as sequences of tokens or structural graph Allamanis et al.
+(2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated program classification Puri et al.
+(2021) and code clone detection Wang et al.
+(2020).
+Existing methods Allamanis et al.
+(2018) for code representation can be roughly divided into two groups, sequence-based representation and graph-based representation.
+Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens Alon et al.
+(2019a).
 Graph-based representation learns the structural information of code.
-Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information (Dinella et al.
-2020; Yasunaga and Liang 2020).
-Unlike code representation, our work focuses on code data preprocessing and studies the powerful pre-trained code models CodeBERT, GraphCodeBERT, and CodeT5 that achieved impressive results on code understanding tasks.
-Data Augmentation for Code Understanding
+Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information Dinella et al.
+(2020); Yasunaga and Liang (2020).
+In this work, we focus on code data preprocessing and study the powerful pre-trained code models and code-specific LLMs that achieved impressive results on code understanding tasks.
+2.2 Data Augmentation for Code Understanding
 Data augmentation is a well-studied technique in the machine learning field that increases the size and diversity of training data by only modifying the existing labeled data, thus, without further human effort.
-For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym (Feng et al.
-2021).
+For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym Feng et al.
+(2021).
 Recently, data augmentation has gained attention from software engineering researchers.
-Various data augmentation methods designed for source code have been proposed (Gao et al.
-2020; Yefet, Alon, and Yahav 2020; Allamanis, Jackson-Flux, and Brockschmidt 2021; Pour et al.
-2021; Bui, Yu, and Jiang 2021; Wang et al.
-2022; Park, Kim, and Han 2023; Yang et al.
-2023a), with the most general framework involving the use of code refactoring, e.g., local variable renaming and dead store, to create more code data without breaking the semantics of the original codes.
-Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information (Lacerda et al.
-2020).
-Existing works (Bielik and Vechev 2020; Yu, Wang, and Wang 2022; de Paula Rodrigues, Braga, and Dahab 2023; Dong et al.
-2023b, 2024) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models.
-For instance, a recent empirical study (Dong et al.
-2024) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation.
-Besides, the SOTA code data augmentation method, MixCode, which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring.
+Various data augmentation methods designed for source code have been proposed Gao et al.
+(2020); Yefet et al.
+(2020); Allamanis et al.
+(2021); Pour et al.
+(2021); Bui et al.
+(2021); Wang et al.
+(2022); Park et al.
+(2023); Yang et al.
+(2023a); Dong et al.
+(2024a), with the most general framework involving the use of code refactoring, e.g., local variable renaming and dead store, to create more code data without breaking the semantics of the original codes.
+Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information Lacerda et al.
+(2020).
+Existing works Bielik and Vechev (2020); Yu et al.
+(2022); de Paula Rodrigues et al.
+(2023); Dong et al.
+(2025, 2024b) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models.
+For instance, a recent empirical study Dong et al.
+(2024b) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation.
+Besides, the SOTA code augmentation method, MixCode, which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring.
 However, as a variant of Mixup specifically designed for code data, MixCode is limited to applications in code classification tasks, as it necessitates the label with one-hot encoding.
-Compared to existing works, we propose a simple yet effective code augmentation framework that is not only unaffected by the dataset and model constraints but also applicable to both classification and non-classification tasks.
-Definition
+However, existing studies Bielik and Vechev (2020); Yu et al.
+(2022) have stated that these code refactoring methods have limited advantages in boosting the training of code models.
+To solve this problem, we propose a simple yet effective code augmentation framework that is not only unaffected by the dataset and model constraints but also applicable to both classification and non-classification tasks.
+2.3 Robustness Testing of Code Model
+Robustness is an important property of DL models which indicates how the DL model can handle data with different distribution of the training data.
+Normally, there are two types of robustness of DL models, natural robustness and adversarial robustness.
+Natural robustness indicates how models perform when facing the natural distribution shift, for example, data in different environments and newly collected data.
+Adversarial robustness refers to the ability of models to defend the adversarial attacks.
+Robustness testing of code models has been addressed recently.
+Researchers proposed distribution shifted code (e.g., codes have the same functionality but are written by different programmers) data to study the natural robustness of code models.
+Besides, multiple adversarial attack methods have been proposed to test the adversarial robustness of code models, for example, the recent work Yang et al.
+(2022) proposed to change the variable name of code to force the model to make wrong predictions.
+In this work, we employ adversarial attack methods Dong et al.
+(2025) proposed for code to evaluate the adversarial robustness of our trained models.
+Following the existing work Dong et al.
+(2023), we construct the natural robustness datasets to evaluate the performance of all code models.
+3 Definition
 We first define the problem of data augmentation for code models.
-Definition 1.
+Definition 1 (Program Code)
 A program code  [^5] For simplicity, we use code in the paper.
-is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: $content$ contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
-Definition 2.
-A code model is defined as a 3-tuple $M=\left(parameter,\mathcal{P},Y\right)$, where $parameter$ contains all the weights in code model $M$, $\mathcal{P}$ is the input code, and $Y$ is the label of $\mathcal{P}$ given a specific downstream code task.
-Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 (Wang et al. 2021).
-Definition 3.
-Given a code $P$, a code transformation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)\&\mathcal{P^{\prime}}.content\neq\mathcal{P}.content$.
-Here, code transformation methods can be any method that can change the content of program statements such as text data transformation methods (e.g., synonym replacement and code refactoring methods (e.g., dead if statement adding).
-Definition 4.
-Given a code model $M$, a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, semantic-preserving (breaking) code augmentation indicates $\mathcal{P}.se=\mathcal{P^{\prime}}.se$ $M(\mathcal{P})=M(\mathcal{P^{\prime}})$ ($\mathcal{P}.se\neq\mathcal{P^{\prime}}.se$ $M(\mathcal{P})\neq M(\mathcal{P^{\prime}})$).
-Definition 5.
+is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: contentcontent contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
+Definition 2 (Code Model)
+A code model is defined as a 3-tuple $M=\left(parameter,\mathcal{P},Y\right)$, where parameterparameter contains all the weights in code model $M$, $\mathcal{P}$ is the input code, and $Y$ is the label of $\mathcal{P}$ given a specific downstream code task.
+Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 Wang et al. (2021).
+Definition 3 (Code Augmentation)
+Given a code $P$, a code augmentation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)\&\mathcal{P^{\prime}}.content\neq\mathcal{P}.content$.
+Here, code augmentation methods can be any method that can change the content of program statements such as text-oriented data augmentation methods (e.g., synonym replacement and code refactoring methods (e.g., dead if statement adding).
+Definition 4 (Semantic-Preserving (Breaking))
+Given a code model $M$, a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, semantic-preserving (breaking) code augmentation indicates $\mathcal{P}.se=\mathcal{P^{\prime}}.se$ & $M(\mathcal{P})=M(\mathcal{P^{\prime}})$ ($\mathcal{P}.se\neq\mathcal{P^{\prime}}.se$ & $M(\mathcal{P})\neq M(\mathcal{P^{\prime}})$).
+Definition 5 (Syntax-Preserving (Breaking))
 Given a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, syntax-preserving (breaking) code augmentation indicates $\mathcal{P}.sy=\mathcal{P^{\prime}}.sy$ ($\mathcal{P}.sy\neq\mathcal{P^{\prime}}.sy$).
-Assumption and Preliminary Study
-GenCode selects the most valuable data for the code model training.
-This approach is grounded on the assumption that the data that code models have less confidence in is more useful for model training.
+4 Assumption and Preliminary Study
+GenCode utilizes the influence score to select the most valuable data for the code model training, where the influence score is defined as the loss value.
+Therefore, this approach is grounded on the assumption that the data that code models have less confidence in is more useful for model training.
 To check whether this assumption is valid or not, we conducted a preliminary study to explore if there is a correlation between the accuracy of the trained code model and the confidence of the model on the training data.
-Specifically, the study is conducted on the Google Code Jam (GCJ) (Yang et al.
-2022) dataset and CodeBERT pre-trained code model.
+Specifically, the study is conducted on the Google Code Jam (GCJ) Yang et al.
+(2022) dataset and CodeBERT pre-trained code model.
 The detailed steps are as follows:
@@
 We then test each code model using the original test data and save 100 accuracies.
-Step 4. We use the Pearson Correlation Coefficient (PCC) (Schober, Boer, and Schwarte 2018) metric to check the correlation between the loss values and the accuracies.
-Table 1: Correlation between loss values and code model accuracy. The P-value less than 0.01 is highlighted with a blue background.
-Coefficient P-value
-Initialized (epoch 0) 0.6075 7.9687e-12
-Middle (epoch 10) 0.2440 0.0125
-Well-trained (epoch 20) 0.1657 0.0926
-Table 1 summarizes the results of our preliminary study.
-Here, we considered three types of code models, the initialed model, the model that has been trained for a few epochs but not converged, and the well-trained code model, to check if our assumption can stand in different scenarios.
+Step 4. We use the Pearson Correlation Coefficient (PCC) Schober et al. (2018) metric to check the correlation between the loss values and the accuracies.
+We consider three training stages, which are the Early stage (Initialized, Epochs 0–9), the Mid stage (Middle, Epochs 10–19), and the Late stage (Well-trained, Epochs 20-29).
+Fig.
+1 visualizes the results of our preliminary study.
+Here, we considered three types of code models, the initialized model, the model that has been trained for a few epochs but not converged, and the well-trained code model, to check if our assumption can stand in different scenarios.
+(a) Initialized
+(b) Middle
+(c) Well-trained
+Figure 1: Correlation between loss values and code model accuracy.
 The results indicate that there is a positive correlation between the loss values of data and the accuracy of code models trained by using these data.
 From the initialized model to the well-trained model, this correlation becomes weaker but still exists.
-We conjecture that when the model is in the early stage, most of the data that have high loss values are useful for the model training.
+We conjecture that when the model is in the early stage, most of the data that have high loss values with more informative and lead to better model generalization.
 However, when training proceeds, the data that still have high loss values are the model itself cannot learn.
-Thus, there will be a correlation drop from the early epoch to the later ones.
-Based on this phenomenon, we can propose an effective and reliable data augmentation framework to boost the training of code models based on this phenomenon.
+As training progresses, the correlation between loss values and accuracy gradually weakens, which may be attributed to overfitting or the model’s limited capacity to learn from inherently difficult samples.
+This observation supports the hypothesis that the strength of this correlation is indicative of the model’s generalization ability.
+In particular, the decline in correlation from early to later training stages highlights the shifting informativeness of high-loss samples.
+Leveraging this insight, we design a data augmentation framework that prioritizes informative samples, especially in the early stages, to enhance the training effectiveness of code models.
 Finding: There is a positive correlation between the loss values of code data and the accuracy of code models trained by using these data.
 When the code model is randomly initialized, this correlation is significant (with a P-value less than 0.01).
-Methodology
-Overview
+5 Methodology
+5.1 Overview
 The goal of GenCode is to produce code models with better performance on test data by training models on carefully generated data.
-Given a training set of program $\mathcal{P}_{train}$, its transformed version $\mathcal{P^{\prime}}_{train}$, and a set of test programs $\mathcal{P}_{test}$, after the training process $\mathsf{train}\left(\cdot\right)$ , $\varrho(M^{\prime}(\mathcal{P}_{test}))>\varrho(M(\mathcal{P}_{test}))$, where $M.parameter=\mathsf{train}(M,\mathcal{P}_{train})$ $M^{\prime}.parameter=\mathsf{train}(M^{\prime},\mathcal{P^{\prime}}_{train})$, and $\varrho\left(\cdot\right)$ is the performance measurement.
-GenCode follows the generation-and-selection workflow and utilizes different types of code transformation methods to generate diverse code data first, then selects code data that are most useful for training as the training data.
+Given a training set of program $\mathcal{P}_{train}$, its transformed version $\mathcal{P^{\prime}}_{train}$, and a set of test programs $\mathcal{P}_{test}$, after the training process $\mathsf{train}\left(\cdot\right)$ , $\varrho(M^{\prime}(\mathcal{P}_{test}))>\varrho(M(\mathcal{P}_{test}))$, where $M.parameter=\mathsf{train}(M,\mathcal{P}_{train})$ & $M^{\prime}.parameter=\mathsf{train}(M^{\prime},\mathcal{P^{\prime}}_{train})$, and $\varrho\left(\cdot\right)$ is the performance measurement.
+GenCode follows the generation-and-selection workflow and utilizes different types of code augmentation methods to generate diverse code data first, then selects code data that are most useful for training as the training data.
 Algorithm 1 shows how GenCode works.
@@
 Here, the data importance is measured by the loss values produced by the code model which is the most common indicator to check how the model performs on the data.
-Overall, GenCode contains two main steps, utilizing code data augmentation methods to compose the search space and importance-guided data selection.
 Algorithm 1 GenCode
 1:$M$: initialized DNN model
-2:$\mathcal{P},Y$: original code data and labels
-3:$\mathcal{T}=\left\{t\right\}:$ a set of code augmentation methods
+[…]
+[…]
 4:$M$: trained model
 5:for $run\leftarrow\{0,\ldots,\#epochs\}$ do
-6:     $\mathcal{P^{\prime}},Y^{\prime}\ \leftarrow\phi,\phi$triangleright initialization
+6:  $\mathcal{P^{\prime}},Y^{\prime}\ \leftarrow\phi,\phi$ $\triangleright$ initialization
 7:  for $t\in\mathcal{T}$ do
 8:   $\mathcal{P^{\prime}}\leftarrow\mathcal{P^{\prime}}\cup t\left(\mathcal{P}\right)$
-9:         $Y^{\prime}\leftarrow Y^{\prime}\cup Y$triangleright code augmentation
-10:     $I\leftarrow\mathsf{Importance}\left(M,\mathcal{P^{\prime}}\right)$triangleright computing importance
-11:     $Index\leftarrow\mathsf{argSort}\left(I\right)\left[:||\mathcal{P}||\right]$triangleright sort by importance
-12:     $X_{train}\leftarrow\mathcal{P^{\prime}}\left[Indext\right]$
-13:     $Y_{train}\leftarrow Y^{\prime}\left[Indext\right]$
+9:   $Y^{\prime}\leftarrow Y^{\prime}\cup Y$ $\triangleright$ code augmentation
+10:  $I\leftarrow\mathsf{Importance}\left(M,\mathcal{P^{\prime}}\right)$ $\triangleright$ computing importance
+11:  $Index\leftarrow\mathsf{argSort}\left(I\right)\left[:||\mathcal{P}||\right]$ $\triangleright$ sort by importance
+12:  $X_{train}\leftarrow\mathcal{P^{\prime}}\left[Index\right]$
+13:  $Y_{train}\leftarrow Y^{\prime}\left[Index\right]$
 14:  $M\leftarrow\mathsf{Train}\left(M,X_{train},Y_{train}\right)$
 15:return $M$
-Search Space
-Even though the code has strict syntax constraints, recent work (Dong et al.
-2023b) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training.
-To enhance the diversity of source code within the feature space, therefore, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 4 and Definition 5).
-GenCode adapts code refactoring methods as semantic/syntax-preserving methods.
-Code refactoring methods follow the constraints of programming languages to edit the code while keeping the main functionality of the code.
-On the other hand, text augmentation methods are used as syntax-breaking methods in GenCode which are widely used in the NLP field.
-For example, text augmentation methods can randomly delete some tokens in the code.
+More specifically, Fig.
+2 depicts the workflow of GenCode in one training epoch.
+First, a series of code augmentation methods are applied to the original dataset to generate an augmented dataset (search space).
+Then, samples are selected based on their influence score, i.e., the loss during model training (importance-guided data selection), ensuring the selected dataset keeps the same size as the original dataset for a fair comparison.
+Figure 2: Workflow of GenCode in one training epoch.
+Overall, GenCode contains two main steps, utilizing code augmentation methods to compose the search space and importance-guided data selection (will be introduced in Section 5.2 and Section 5.3, respectively).
+5.2 Search Space
+Even though the code has strict syntax constraints, recent work Dong et al.
+(2025) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training.
+To enhance the diversity of source code within the feature space, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 4 and Definition 5).
+For instance, synonym replacement (SR) may turn an identifier ”count” into a natural-language word, such as ”amount”, which is not declared in the symbol table and thus results in a compilation error.
+In contrast, refactoring-based renaming updates all references consistently and guarantees syntactic correctness.
+GenCode adapts code refactoring methods (i.e., API renaming and Duplication) as semantic/syntax-preserving methods.
+Code refactoring methods follow the constraints of programming languages (PL) to edit the code while keeping the main functionality of the code.
+On the other hand, text-oriented data augmentation methods Dong et al.
+(2025) are used as syntax-breaking methods in GenCode, which are widely used in the NLP field.
+For example, text-oriented data augmentation methods can randomly delete some tokens in the code.
 However, even though the syntax has been broken, the code model can still successfully predict those codes since it reads the code as a sequence.
-In total, as shown in Table 2, GenCode currently supports 18 code refactoring methods and 5 text augmentation methods.
-Besides, Figure 1 provides an example of the training candidate that is to generate the search space.
+In total, as shown in Table 1, GenCode currently supports 18 code refactoring methods and 5 text-oriented data augmentation methods.
+Besides, Fig.
+3 provides an example of the training candidate that is to generate the search space.
 In the syntax-preserving, we present a code refactoring method named API Renaming, which involves modifying the function name from ”main()” to ”even()”.
-In the syntax-breaking, we demonstrate another instance of API Renaming, where two different statements within a program are randomly selected and their positions swapped.
+In the syntax-breaking, we demonstrate another instance of Random Swap, where two different statements within a program are randomly selected and their positions swapped.
 As depicted in this example, the third and fourth lines have been interchanged.
@@
 API renaming
-Filed enhancement
-Synonym replacement (SR) Random insertion (RI) Random swap (RS Random deletion (RD) Back-translation (BT)
+Field enhancement
+Synonym replacement (SR) Random insertion (RI) Random swap (RS) Random deletion (RD) Back-translation (BT)
 Arguments adding
@@
 Figure 3: An example of atomic data augmentation operators.
-Table 3: Effectiveness of data augmentation methods w.r.t.
-test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data and ASR $\downarrow$ (%) on test data.
-The best results are highlighted in blue.
-Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
-GCJ Refactory Java250 BigCloneBench
-Model DA method Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$
-No Aug 90.96 ± 0.07 33.31 56.13 96.21 ± 0.13 36.31 42.31 96.36 ± 0.09 40.55 53.37 96.91 ± 0.12 9.74 24.61
-Refactor 91.74 ± 0.11 35.56 52.84 95.95 ± 0.19 32.64 40.89 96.44 ± 0.12 38.35 52.56 96.97 ± 0.21 8.55 18.29
-SR 70.66 ± 0.11 35.46 52.71 96.71 ± 0.11 39.25 53.91 96.21 ± 0.02 42.01 53.79 97.02 ± 0.15 12.52 26.11
-RI 59.91 ± 0.21 40.78 65.76 96.43 ± 0.18 38.45 45.57 96.15 ± 0.06 43.15 55.07 97.38 ± 0.23 8.31 30.08
-RS 93.24 ± 0.13 28.26 52.44 96.73 ± 0.15 38.63 45.98 96.49 ± 0.04 39.36 50.79 96.92 ± 0.12 13.41 30.13
-RD 76.98 ± 0.09 35.65 55.46 96.46 ± 0.18 30.23 41.32 96.61 ± 0.12 38.31 52.42 97.11 ± 0.19 9.58 30.15
-BT 82.74 ± 0.15 40.38 55.95 94.31 ± 0.38 62.98 69.87 96.06 ± 0.06 42.21 53.32 97.13 ± 0.17 11.29 26.32
-MixCode 92.21 ± 0.25 29.28 52.53 96.76 ± 0.31 31.04 41.01 96.62 ± 0.23 36.14 50.67 - - -
-CodeBERT GenCode 96.23 ± 0.18 22.33 46.62 99.32 ± 0.14 23.75 31.25 98.89 ± 0.09 29.89 40.65 98.93 ± 0.11 7.12 15.83
-No Aug 93.96 ± 0.15 31.23 56.02 96.84 ± 0.16 22.97 30.26 96.41 ± 0.11 23.16 42.66 96.88 ± 0.09 4.19 12.22
-Refactor 91.72 ± 0.21 30.31 56.58 95.49 ± 0.21 24.26 32.57 96.53 ± 0.15 23.13 42.42 97.11 ± 0.23 6.05 6.17
-SR 89.46 ± 0.13 40.02 66.68 95.57 ± 0.13 28.39 47.15 96.46 ± 0.14 22.92 42.18 97.14 ± 0.16 6.13 10.16
-RI 80.47 ± 0.21 43.51 65.75 94.51 ± 0.25 25.08 32.87 96.52 ± 0.09 22.37 42.33 97.35 ± 0.17 2.23 6.28
-RS 94.76 ± 0.15 30.18 54.75 97.88 ± 0.15 21.85 29.67 96.44 ± 0.14 22.68 41.98 97.54 ± 0.12 6.03 11.94
-RD 91.71 ± 0.26 31.73 60.92 96.53 ± 0.19 20.35 26.98 96.56 ± 0.21 22.39 41.84 97.16 ± 0.21 6.18 8.27
-BT 80.47 ± 0.05 41.65 62.99 95.99 ± 0.27 35.56 46.23 96.39 ± 0.13 22.91 42.32 97.11 ± 0.16 8.18 14.12
-MixCode 94.19 ± 0.25 30.26 54.49 97.29 ± 0.15 14.89 26.53 96.89 ± 0.19 20.03 42.01 - - -
-GraphCodeBERT GenCode 97.93 ± 0.14 26.43 48.87 99.45 ± 0.17 15.78 21.92 99.16 ± 0.13 18.76 33.81 99.23 ± 0.08 1.86 4.62
-No Aug 95.02 ± 0.07 30.78 55.67 97.22 ± 0.17 22.45 29.69 96.98 ± 0.08 31.76 49.82 97.38 ± 0.13 10.32 22.97
-Refactor 92.21 ± 0.14 30.12 55.93 96.43 ± 0.22 28.76 35.45 96.92 ± 0.11 30.66 47.12 97.24 ± 0.12 9.32 22.05
-SR 90.14 ± 0.21 39.21 51.84 94.67 ± 0.14 28.24 36.26 96.53 ± 0.05 28.65 46.01 97.31 ± 0.08 15.87 25.98
-RI 78.77 ± 0.05 41.02 63.46 95.02 ± 0.15 24.78 33.61 96.11 ± 0.09 27.14 46.76 97.46 ± 0.13 10.43 29.76
-RS 95.66 ± 0.14 29.35 51.74 97.97 ± 0.29 21.32 27.76 96.34 ± 0.11 28.45 48.45 97.03 ± 0.15 14.33 30.43
-RD 92.01 ± 0.11 31.04 56.85 97.52 ± 0.26 19.93 24.87 97.21 ± 0.07 26.09 45.24 97.43 ± 0.07 9.98 28.87
-BT 81.45 ± 0.13 39.65 54.24 95.38 ± 0.11 33.62 43.67 96.45 ± 0.14 29.47 45.38 97.65 ± 0.14 10.43 25.34
-MixCode 94.32 ± 0.16 29.82 42.56 97.32 ± 0.13 12.49 18.32 96.91 ± 0.11 28.73 48.87 - - -
-CodeT5 GenCode 98.85 ± 0.11 23.57 44.76 99.64 ± 0.06 12.13 19.56 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
-Importance-Guided Data Selection
+5.3 Importance-Guided Data Selection
 After constructing the training candidate, we then choose the most valuable code data that can benefit the code model training.
@@
 That means, GenCode takes all the transformed code as input to the code model and records all the loss values first, then ranks all the code in descending order based on the loss values and selects the first $K$ code as the training data.
-Following the common setting in the data augmentation problem (Dong et al.
-2023a), $K$ is set as the number of original training data.
+Following the common setting in the data augmentation problem Dong et al.
+(2023), $K$ is set as the number of original training data.
 GenCode chooses loss value as the indicator since it can be generalized to any code model and task.
-Evaluation
-We evaluate both the accuracy and robustness of trained code models. Besides, we check how does importance score affects the effectiveness of GenCode.
-Datasets and Models
-Datasets and Models We evaluate GenCode on four code understanding tasks, bug detection (Hu et al.
-2019), authorship attribution (Yang et al.
-2022), code clone detection (Yang et al.
-2022) (Note that, MixCode is not suitable for this task since its input is code pairs.), and problem classification (Puri et al.
-2021).
+6 Evaluation Setup
+We evaluate both the accuracy and robustness of trained code models. Besides, we check how does influence score affects the effectiveness of GenCode.
+6.1 Study Design
+We conduct experiments to evaluate the usefulness of GenCode and answer the following research questions.
+RQ1: How effective is GenCode in producing accurate code models?
+Accuracy is the fundamental metric that is used to evaluate the performance of a trained model.
+Therefore, we first study the basic DNN property, the test accuracy, of each selected pre-trained code model.
+Specifically, we compare GenCode with models trained without using data augmentation and models trained by baseline data augmentation methods (will be introduced in Section 6.4), and we additionally provide the evaluation results of text-oriented data augmentation methods (as mentioned in Section 5.2) to demonstrate the usefulness of GenCode.
+Additionally, to ensure a fair comparison in terms of training cost, we apply each data augmentation method to every sample in the original training dataset, resulting in an augmented dataset of the same size.
+These augmented datasets are fixed prior to training and remain unchanged across all training epochs.
+RQ2: How effective is GenCode in producing robust code models?
+Robustness is another important property of DNN models.
+In this study, we mainly focus on adversarial robustness and natural robustness.
+Specifically, for the adversarial robustness, we utilize two widely used adversarial attack methods including MHM Zhang et al.
+(2020) and ALERT Yang et al.
+(2022), to assess the performance of code models trained by the data augmentation method in robustness improvement.
+For natural robustness, we randomly apply a single code refactoring method Dong et al.
+(2023) to each sample in order to construct a transformed code dataset that preserves the original semantic labels, thereby enabling a comprehensive evaluation of the generalization capability under natural variations.
+RQ3: How does the influence score affect the effectiveness of GenCode?
+The importance of score-based data selection is the key component is GenCode.
+Therefore, in this RQ, we change the selection strategies, i.e., selecting data with minimum confidence and randomly selecting data, to study the contribution of our proposed selection strategy.
+To achieve this, we repeat the experiments outlined in RQ1 and RQ2.
+Subsequently, we investigate how the selection of different scores influences the performance of pre-trained models.
+Note that, for a fair comparison, we evaluate only the adversarial robustness, as the use of random may artificially inflate performance on the transformed dataset Dong et al.
+(2023).
+RQ4: How does the data selection method affect the effectiveness of GenCode?
+Data selection strategies play a critical role in determining the performance of trained code models Hu et al.
+(2024).
+In this RQ, we select representative methods from the literature to serve as baselines for comparison with GenCode.
+Specifically, we include LIME Ribeiro et al.
+(2016), which provides interpretable, local approximations of complex model predictions and can be leveraged during the model training phase to inform data selection by identifying influential samples.
+As one of the classical data selection methods, Influence Function Koh and Liang (2017) estimates the effect of individual training samples by approximating how infinitesimal perturbations to the training data impact the learned parameters, and can be effectively utilized during the phase of model training to guide data selection.
+Additionally, we incorporate BADGE Ash et al.
+(2019), which selects data by applying K-means clustering on gradient embeddings, and has been shown to be a stable and effective active learning method in recent empirical studiesHu et al.
+(2024).
+We evaluate the performance of selected data selection methods in terms of both accuracy and natural robustness.
+Adversarial robustness is not considered in this study, as existing adversarial attacks in RQ2 are not well-suited for LLMs.
+6.2 Datasets and Models
+We evaluate GenCode on four code understanding tasks as shown in Table 2, bug detection Hu et al.
+(2019), authorship attribution Yang et al.
+(2022), code clone detection Yang et al.
+(2022) (Note that, MixCode is not suitable for this task since its input is code pairs.), and problem classification Puri et al.
+(2021).
+We conduct all experiments on a server equipped with 4 NVIDIA RTX A6000 GPUs.
 For each code task, we fine-tune three types of models based on pre-trained code models, CodeBERT, GraphCodeBERT, and CodeT5.
-For the model training, we follow existing works (Hu et al.
-2019; Wang et al.
-2021; Puri et al.
-2021; Yang et al.
-2022; Dong et al.
-2023a), to set all configurations.
+For the model training, we follow existing works Hu et al.
+(2019); Wang et al.
+(2021); Puri et al.
+(2021); Yang et al.
+(2022); Dong et al.
+(2023), to set all configurations.
 Specifically, we set the training epoch as 50 for all tasks and models.
-Regarding the optimizer, we utilize Adam (Kingma and Ba 2014) with a learning rate of $10^{-3}$ for all the models.
+Regarding the optimizer, we utilize Adam Kingma and Ba (2014) with a learning rate of $10^{-3}$ for all the models.
 Besides, we use early stopping with a patience of 20 to mitigate overfitting.
@@
 For MixCode, we follow the recommendation of the original paper and set the Mixup ratio as $\alpha=0.1$.
-For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  (Yang et al.
-2022).
-Evaluation Metrics.
+For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  Yang et al.
+(2022).
+Considering the code-specific Large Language Models (LLMs), we select StarCoder2 Lozhkov et al.
+(2024), a representative model from the BigCode  [^6] https://www.bigcode-project.org project, which is a collaborative initiative dedicated to the responsible development of LLMs for Code, alongside the recently released Qwen2.5-Coder Hui et al.
+(2024).
+Table 2: Details of tasks, datasets, and DNN models. #Training and #Test are the number of data for training and testing, respectively.
+Dataset Language Task #Training #Test Model
+Google Code Jam (GCJ) Python Authorship attribution 528 132 CodeBERT GraphCodeBERT CodeT5 StarCoder2-7B Qwen2.5-Coder-7B
+Refactory Python Bug detection 3,380 423
+Java250 Java Problem classification 48,000 15,000 CodeBERT GraphCodeBERT CodeT5 StarCoder2-7B Qwen2.5-Coder-7B
+BigCloneBench Java Clone detection 90,102 4,000
+6.3 Configuration
+We conduct all experiments on a computation server equipped with 4× NVIDIA A100 (40GB) GPUs, two Intel Xeon Gold 6330 CPUs, and 512GB RAM.
+To ensure fairness and consistency across benchmarks, we adopt a unified training configuration following established practices in prior work Wang et al.
+(2021); Dong et al.
+(2023); Hu et al.
+(2019); Yang et al.
+(2022); Puri et al.
+(2021).
+Specifically, we train CodeBERT, GraphCodeBERT, and CodeT5 for 50 epochs, and train StarCoder2-7B and Qwen2.5-Coder-7B for 20 epochs, using the Adam optimizer (learning rate = 1e-3, weight decay = 0.01).
+The batch size is fixed to 32, and the maximum sequence length is set to 512 tokens for all encoder-based architectures.
+To mitigate overfitting and ensure stable convergence, we apply early stopping with a patience of 20 epochs.
+Each experiment is executed five independent times, and we report the mean and standard deviation to account for stochastic variability.
+For MixCode, we follow the settings recommended in the original paper and adopt a mixup coefficient of $\alpha=0.1$.
+For adversarial training baselines, we directly use the default hyperparameters released by their authors to avoid unintended deviations from the original configurations.
+To support reproducibility, we provide the complete training pipeline 4, including data preprocessing scripts, model configuration files (e.g., batch size, sequence length, and optimizer parameters), and execution instructions.
+This ensures that all experimental results presented in this work can be reproduced under the same computational environment and hyperparameter settings.
+6.4 Evaluation metrics and Baselines
 In the evaluation, we assess the performance of trained models from two perspectives, namely, clean accuracy and model robustness.
-For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT (Yang et al.
-2022), which uses variable renaming while considering natural semantics, and MHM (Zhang et al.
-2020), which also randomly replaces variable names but does not account for natural semantics.
-We do not consider the natural robustness since the existing natural robustness datasets can not fully support our studied tasks, i.e., no natural robustness datasets support authorship attribution, bug detection, and code clone detection.
-Baseline Methods.
-We consider model training without data augmentation (No Aug), Refactor (Yu, Wang, and Wang 2022), a family of text-oriented data augmentation methods (Dong et al.
-2023b), and MixCode as baseline methods.
+For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT Yang et al.
+(2022), which uses variable renaming while considering natural semantics, and MHM Zhang et al.
+(2020), which also randomly replaces variable names but does not account for natural semantics.
+Metric ASR evaluates the performance of models, including CodeBERT, GraphCodeBERT, and CodeT5 Dong et al.
+(2025).
+In addition, we evaluate the natural robustness of all trained models by applying 18 different code refactoring methods to generate diverse program variants.
+Specifically, following the existing work Dong et al.
+(2025, 2023), we construct transformed versions of each code sample in the dataset by randomly applying one of the refactoring methods.
+We consider model training without data augmentation (No Aug), Refactor Yu et al.
+(2022), a family of text-oriented data augmentation methods Dong et al.
+(2025), and MixCode as baseline methods.
 Refactor, which keeps code syntax rules, is a widely used data augmentation framework specifically designed for code data.
-In contrast, text-oriented code data augmentation methods, which slightly break syntax rules, have recently proven more effective than Refactor.
-Additionally, MixCode is a more recent and high-performing code data augmentation method that significantly enhances code models.
-Accuracy Evaluation
+In contrast, text-oriented data augmentation methods, which slightly break syntax rules, have recently proven more effective than Refactor.
+Additionally, MixCode is a more recent and high-performing code augmentation method that significantly enhances code models.
+7 Results Analysis
+7.1 RQ1: How effective is GenCode in producing accurate code models?
 Overall accuracy.
-Column Accuracy in Table 3 summarizes the final test accuracy of code models trained by using different code data augmentation methods including GenCode.
+Column Accuracy in Table 3 summarizes the final test accuracy of code models trained by using different code augmentation methods including GenCode.
 Surprisingly, from the results, we can see that most of the existing code augmentation methods except RS and MixCode negatively affect the performance of code models.
-This finding is similar to the conclusion from recent code augmentation studies (Bielik and Vechev 2020; Yu, Wang, and Wang 2022; Dong et al.
-2023b) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
+This finding is similar to the conclusion from recent code augmentation studies Bielik and Vechev (2020); Yu et al.
+(2022); Dong et al.
+(2025) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
+Table 3: Effectiveness of data augmentation methods w.r.t.
+test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data.
+No Aug: without data augmentation.
+The best results are highlighted in blue.
+Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
+Model DA method GCJ Refactory Java250 BigCloneBench
+No Aug 90.96 ± 0.07 96.21 ± 0.13 96.36 ± 0.09 96.91 ± 0.12
+Refactor 91.74 ± 0.11 95.95 ± 0.19 96.44 ± 0.12 96.97 ± 0.21
+SR 70.66 ± 0.11 96.71 ± 0.11 96.21± 0.02 97.02 ± 0.15
+RI 59.91 ± 0.21 96.43 ± 0.18 96.15 ± 0.06 97.38 ± 0.23
+RS 93.24 ± 0.13 96.73 ± 0.15 96.49 ± 0.04 96.92 ± 0.12
+RD 76.98 ± 0.09 96.46 ± 0.18 96.61 ± 0.12 97.11 ± 0.19
+BT 82.74 ± 0.15 94.31 ± 0.38 96.06 ± 0.06 97.13 ± 0.17
+MixCode 92.21 ± 0.25 96.76 ± 0.31 96.62 ± 0.23 -
+CodeBERT GenCode 96.23 ± 0.18 99.32 ± 0.14 98.89 ± 0.09 98.93 ± 0.11
+No Aug 93.96 ± 0.15 96.84 ± 0.16 96.41 ± 0.11 96.88 ± 0.09
+Refactor 91.72 ± 0.21 95.49 ± 0.21 96.53 ± 0.15 97.11 ± 0.23
+SR 89.46 ± 0.13 95.57 ± 0.13 96.46 ± 0.14 97.14 ± 0.16
+RI 80.47 ± 0.21 94.51 ± 0.25 96.52 ± 0.09 97.35 ± 0.17
+RS 94.76 ± 0.15 97.88 ± 0.15 96.44 ± 0.14 97.54 ± 0.12
+RD 91.71 ± 0.26 96.53 ± 0.19 96.56 ± 0.21 97.16 ± 0.21
+BT 80.47 ± 0.05 95.99 ± 0.27 96.39 ± 0.13 97.11 ± 0.16
+MixCode 94.19 ± 0.25 97.29 ± 0.15 96.89 ± 0.19 -
+GraphCodeBERT GenCode 97.93 ± 0.14 99.45 ± 0.17 99.16 ± 0.13 99.23 ± 0.08
+No Aug 95.02 ± 0.07 97.22 ± 0.17 96.98 ± 0.08 97.38 ± 0.13
+Refactor 92.21 ± 0.14 96.43 ± 0.22 96.92 ± 0.11 97.24 ± 0.12
+SR 90.14 ± 0.21 94.67 ± 0.14 96.53 ± 0.05 97.31 ± 0.08
+RI 78.77 ± 0.05 95.02 ± 0.15 96.11 ± 0.09 97.46 ± 0.13
+RS 95.66 ± 0.14 97.97 ± 0.29 96.34 ± 0.11 97.03 ± 0.15
+RD 92.01 ± 0.11 97.52 ± 0.26 97.21 ± 0.07 97.43 ± 0.07
+BT 81.45 ± 0.13 95.38 ± 0.11 96.45 ± 0.14 97.65 ± 0.14
+MixCode 94.32 ± 0.16 97.32 ± 0.13 96.91 ± 0.11 -
+CodeT5 GenCode 98.85 ± 0.11 99.64 ± 0.06 99.38 ± 0.11 99.34 ± 0.12
+No Aug 98.91 ± 0.02 98.12 ± 0.01 97.83 ± 0.02 98.75 ± 0.03
+Refactor 96.92 ± 0.01 97.99 ± 0.03 97.25 ± 0.02 98.61 ± 0.01
+SR 92.36 ± 0.04 97.27 ± 0.04 97.27 ± 0.03 98.46 ± 0.04
+RI 90.47 ± 0.05 96.82 ± 0.03 97.26 ± 0.02 98.31 ± 0.02
+RS 97.66 ± 0.06 97.41 ± 0.06 98.19 ± 0.01 98.15 ± 0.01
+RD 98.33 ± 0.03 97.02 ± 0.02 97.72 ± 0.02 98.36 ± 0.03
+BT 89.92 ± 0.02 97.09 ± 0.05 97.41 ± 0.02 98.12 ± 0.05
+MixCode 98.56 ± 0.04 98.21 ± 0.08 98.23 ± 0.03 -
+StarCoder2-7B GenCode 98.93 ± 0.04 99.65 ± 0.02 99.41 ± 0.05 99.36 ± 0.09
+No Aug 98.95 ± 0.03 98.19 ± 0.03 98.11 ± 0.02 98.78 ± 0.01
+Refactor 97.22 ± 0.03 97.51 ± 0.02 97.31 ± 0.01 98.32 ± 0.04
+SR 92.45 ± 0.02 97.47 ± 0.03 97.37 ± 0.04 98.51 ± 0.05
+RI 90.72 ± 0.03 96.31 ± 0.02 98.07 ± 0.05 98.34 ± 0.04
+RS 96.89 ± 0.05 97.52 ± 0.05 97.31 ± 0.02 98.02 ± 0.06
+RD 96.83 ± 0.01 96.89 ± 0.04 97.52 ± 0.03 98.25 ± 0.04
+BT 91.07 ± 0.06 97.15 ± 0.03 98.02 ± 0.02 98.06 ± 0.08
+MixCode 98.61 ± 0.03 98.46 ± 0.06 98.53 ± 0.04 -
+Qwen2.5-Coder-7B GenCode 98.96 ± 0.04 99.65 ± 0.01 99.42 ± 0.04 99.36 ± 0.05
 Then, compared to the existing data augmentation methods, the results demonstrate that GenCode consistently enhances the training and achieves the best results across all the considered datasets and models.
-Concretely, on average, GenCode generates models with 4.52%, 3.05%, 2.56%, and 2.11% higher accuracy than No Aug in GCJ, Refactory, Java250, and BigCloneBench, respectively.
+Concretely, on average, for pre-trained code models, GenCode generates models with 4.52%, 3.05%, 2.56%, and 2.11% higher accuracy than No Aug in GCJ, Refactory, Java250, and BigCloneBench, respectively.
 Besides, compared to MixCode, GenCode can achieve 4.1%, 2.34%, and 2.33% better results in GCJ, Refactory, and Java250 datasets, respectively.
-Statistical analysis.
-We conducted statistical tests, including the T-Test and Wilcoxon signed-rank test, across all pre-trained models.
-The results show that GenCode significantly (with p-value $<$ 0.05) outperforms other methods including MixCode.
-In conclusion, GenCode is effective in training accurate code models.
+Moving to LLMs, GenCode improves the performance of StarCoder2 across all tasks by up to 0.93% on average, and achieves a 0.84% average improvement for Qwen2.5-Coder.
+Convergence speed analysis.
+Fig.
+4 depicts the average test accuracy at each training epoch of CodeBERT using different data augmentation methods.
+The results demonstrated that GenCode consistently outperforms other baseline methods during the whole training process.
+We can conclude that the choice of the number of training epochs will not affect the advantages of our method over other methods.
 (a) Authorship attribution
@@
 (d) Clone detection
-Figure 2: Convergence speed of CodeBERT using different code data augmentation methods in each task.
-Convergence speed analysis.
-Figure 2 depicts the average test accuracy at each training epoch of CodeBERT using different data augmentation methods.
-The results demonstrated that GenCode consistently outperforms other baseline methods during the whole training process.
-We can conclude that the choice of the number of training epochs will not affect the advantages of our method over other methods.
+Figure 4: Convergence speed of CodeBERT using different code augmentation methods in each task.
+Visualizing principal component analysis.
+Furthermore, we explore why models trained by GenCode have higher accuracy than models trained using other methods by analyzing the code representation capability of trained code embeddings.
+Specifically, 1) first, we extract the code embeddings produced by each trained model using all the test data as code features.
+2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) Maćkiewicz and Ratajczak (1993) algorithm.
+3) After that, we visualize the reduced 2-dimensional vectors.
+Normally, the features extracted by a well-trained model are more clearly distributed based on the data labels.
 (a) Refactor
@@
 Figure 5: Visualization of code embeddings after dimension reduction using Principal Component Analysis (PCA). Model: CodeBERT, dataset: Refactory, task: Bug detection.
-Considering Figure 3, we can see that the embedding produced by GenCode is clearly distributed compared to other methods.
-Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Figure 3.
-The results are shown in the following Table.
-The results confirm that GenCode can better distinguish the two classes.
+Fig.
+5 depicts the visualized features extracted from CodeBERT using the Refactory dataset.
+We can see that the code embeddings trained by using GenCode are more clearly distributed than embedding extracted by other models, i.e., the buggy code and correct code are greatly different in the figure.
+These results indicate that GenCode can train good code embeddings and, therefore, produce better accurate code models for code understanding tasks.
+Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Fig.
+5.
+The results from Table 4 confirm that GenCode can better distinguish the two classes.
 Interestingly, we found that MixCode has the smallest distance result.
 That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries.
-This phenomenon is similar to the findings in (Verma et al.
-2019), where the representations obtained by interpolating hidden states have two desirable properties: the class-representations are flattened into a minimal amount of directions of variation, and all points in between these flat representations, most unobserved during training and off the data manifold, are assigned low-confidence predictions.
+This phenomenon is similar to the findings in Verma et al.
+(2019), where interpolating hidden state representations yields two beneficial properties: 1) class representations are flattened into a minimal number of directions of variation, and 2) all points between these flat representations are assigned low-confidence predictions.
 Table 4: Average distance between the data points from different classes. The best is highlighted with a blue background, and the worst is highlighted with a gray background.
@@
 Average distance 7.03 6.00 4.05 8.78
+Table 5: Results of statistical tests on Accuracy & Natural Robustness.
+A gray background highlights that the comparison is statistically significant.
+Statistical tests method: Wilcoxon signed-rank test.
+Across CodeBERT, GraphCodeBERT, CodeT5, StarCoder2, and Qwen2.5-Coder models.
+DA method No Aug Refactor SR RI RS RD BT MixCode
+Accuracy GenCode 1.907e-06 1.907e-06 1.907e-06 8.845e-05 1.907e-06 1.907e-06 1.907e-06 0.001
+Natural Robustness GenCode 1.907e-06 - 1.907e-06 1.907e-06 1.907e-06 1.316e-04 1.907e-06 6.533e-04
+Statistical analysis.
+We conducted statistical tests with Wilcoxon signed-rank test, across all code models.
+The results from the upper of Table 5 show that GenCode significantly (with p-value << 0.05) outperforms other methods including MixCode.
+In conclusion, GenCode is effective in training accurate code models.
+Answer to RQ1: GenCode consistently outperforms baselines in training more accurate code understanding models.
+Specifically, for pre-trained code models, GenCode yields an average accuracy improvement of 3.06% over models trained without data augmentation and 2.92% over those trained with the SOTA MixCode method.
+For LLMs, GenCode achieves average improvements of 0.93% for StarCoder2 and 0.84% for Qwen2.5-Coder.
+7.2 RQ2: How effective is GenCode in producing robust code models?
+Adversarial robustness evaluation.
+Table 6: Effectiveness of data augmentation methods w.r.t.
+ASR $\downarrow$ (%) on test data.
+The best results are highlighted in blue.
+Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
+GCJ Refactory Java250 BigCloneBench
+Model DA method MHM ALERT MHM ALERT MHM ALERT MHM ALERT
+No Aug 33.31 56.13 36.31 42.31 40.55 53.37 9.74 24.61
+Refactor 35.56 52.84 32.64 40.89 38.35 52.56 8.55 18.29
+SR 35.46 52.71 39.25 53.91 42.01 53.79 12.52 26.11
+RI 40.78 65.76 38.45 45.57 43.15 55.07 8.31 30.08
+RS 28.26 52.44 38.63 45.98 39.36 50.79 13.41 30.13
+RD 35.65 55.46 30.23 41.32 38.31 52.42 9.58 30.15
+BT 40.38 55.95 62.98 69.87 42.21 53.32 11.29 26.32
+MixCode 29.28 52.53 31.04 41.01 36.14 50.67 - -
+CodeBERT GenCode 22.33 46.62 23.75 31.25 29.89 40.65 7.12 15.83
+No Aug 31.23 56.02 22.97 30.26 23.16 42.66 4.19 12.22
+Refactor 30.31 56.58 24.26 32.57 23.13 42.42 6.05 6.17
+SR 40.02 66.68 28.39 47.15 22.92 42.18 6.13 10.16
+RI 43.51 65.75 25.08 32.87 22.37 42.33 2.23 6.28
+RS 30.18 54.75 21.85 29.67 22.68 41.98 6.03 11.94
+RD 31.73 60.92 20.35 26.98 22.39 41.84 6.18 8.27
+BT 41.65 62.99 35.56 46.23 22.91 42.32 8.18 14.12
+MixCode 30.26 54.49 14.89 26.53 20.03 42.01 - -
+GraphCodeBERT GenCode 26.43 48.87 15.78 21.92 18.76 33.81 1.86 4.62
+No Aug 30.78 55.67 22.45 29.69 31.76 49.82 10.32 22.97
+Refactor 30.12 55.93 28.76 35.45 30.66 47.12 9.32 22.05
+SR 39.21 51.84 28.24 36.26 28.65 46.01 15.87 25.98
+RI 41.02 63.46 24.78 33.61 27.14 46.76 10.43 29.76
+RS 29.35 51.74 21.32 27.76 28.45 48.45 14.33 30.43
+RD 31.04 56.85 19.93 24.87 26.09 45.24 9.98 28.87
+BT 39.65 54.24 33.62 43.67 29.47 45.38 10.43 25.34
+MixCode 29.82 42.56 12.49 18.32 28.73 48.87 - -
+CodeT5 GenCode 23.57 44.76 12.13 19.56 21.63 39.76 7.39 12.15
+Column MHM and ALERT in Table 6 summarize the attack success rate of adversarial attack methods against code models.
+First, we can see that similar to the observation in Accuracy Evaluation, not all data augmentation methods can improve the robustness of code models.
+This finding is consistent with the existing works Dong et al.
+(2025).
+Specifically, in the baseline methods, only RS and MixCode have a positive impact on the robustness of code models.
+Considering our method, GenCode consistently improves the robustness of code models and achieves the best results in 21 out of 24 cases.
+On average, compared to No Aug, models trained by GenCode have an 8.42% lower ASR (higher robustness).
+This result breaks the previous finding Bielik and Vechev (2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness.
+On the other hand, compared to MixCode, GenCode can train code models with 4.90% higher robustness on average.
+Natural robustness evaluation.
+Table 7 presents the result of the natural robustness evaluation.
+For a fair comparison, we remove Refactor from the column DA method.
+The results indicate that most text-oriented data augmentation methods Dong et al.
+(2025), such as BT and RI, do not contribute significantly to improving natural robustness.
+Sampling-based code augmentation methods, such as MixCode, yield limited improvements.
+For instance, in code multi-classification tasks including authorship attribution (GCJ), MixCode achieves an average improvement of only 0.56% compared to training without data augmentation.
+Although GenCode outperforms all baselines in Table 7, the improvement remains limited, particularly for LLMs.
+For example, GenCode achieves an average improvement of only 0.98% for StarCoder2-7B and 0.84% for Qwen2.5-Coder-7B across all tasks.
+This observation aligns with findings from existing studies Dong et al.
+(2025); Ding et al.
+(2024), suggesting that the effectiveness of data augmentation diminishes as model scale increases.
+Table 7: Effectiveness of data augmentation methods w.r.t.
+test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on transformed test data.
+No Aug: without data augmentation.
+The best results are highlighted in blue.
+Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
+Model DA method GCJ Refactory Java250 BigCloneBench
+No Aug 87.57 ± 0.13 92.12 ± 0.09 89.68 ± 0.16 91.61 ± 0.18
+SR 62.78 ± 0.22 89.87 ± 0.15 85.51 ± 0.07 86.21 ± 0.12
+RI 50.35 ± 0.19 90.23 ± 0.19 85.21 ± 0.11 92.02 ± 0.13
+RS 88.23 ± 0.13 90.56 ± 0.12 89.97 ± 0.14 85.65 ± 0.11
+RD 69.34 ± 0.18 93.78 ± 0.11 90.34 ± 0.13 91.87 ± 0.12
+BT 75.43 ± 0.15 82.35 ± 0.21 85.23 ± 0.15 87.78 ± 0.09
+MixCode 89.11 ± 0.13 95.36 ± 0.09 93.19 ± 0.12 -
+CodeBERT GenCode 91.23 ± 0.14 95.63 ± 0.14 93.62 ± 0.09 92.56 ± 0.15
+No Aug 87.86 ± 0.12 92.04 ± 0.03 90.58 ± 0.08 91.89 ± 0.09
+SR 66.45 ± 0.09 89.12 ± 0.06 91.05 ± 0.12 89.23 ± 0.13
+RI 53.87 ± 0.13 91.29 ± 0.07 90.88 ± 0.06 93.57 ± 0.15
+RS 87.73 ± 0.15 92.56 ± 0.02 91.44 ± 0.11 89.02 ± 0.08
+RD 73.59 ± 0.08 92.87 ± 0.05 92.01 ± 0.05 92.21 ± 0.16
+BT 73.24 ± 0.11 86.05 ± 0.11 93.14 ± 0.04 90.25 ± 0.12
+MixCode 88.13 ± 0.13 95.24 ± 0.12 94.05 ± 0.15 -
+GraphCodeBERT GenCode 89.78 ± 0.16 95.67 ± 0.15 94.89 ± 0.08 93.89 ± 0.11
+No Aug 89.31 ± 0.15 92.89 ± 0.08 91.13 ± 0.12 91.93 ± 0.13
+SR 68.94 ± 0.06 90.21 ± 0.12 88.23 ± 0.08 89.43 ± 0.08
+RI 58.34 ± 0.18 91.14 ± 0.14 91.13 ± 0.13 93.29 ± 0.17
+RS 89.98 ± 0.19 92.89 ± 0.13 90.87 ± 0.07 88.34 ± 0.07
+RD 75.57 ± 0.13 93.98 ± 0.12 91.97 ± 0.11 92.35 ± 0.14
+BT 75.64 ± 0.11 88.79 ± 0.08 94.06 ± 0.16 90.87 ± 0.05
+MixCode 88.45 ± 0.14 95.86 ± 0.14 93.89 ± 0.19 -
+CodeT5 GenCode 91.38 ± 0.09 95.87 ± 0.11 95.06 ± 0.12 94.03 ± 0.12
+No Aug 92.53 ± 0.02 94.12 ± 0.05 92.71 ± 0.11 93.21 ± 0.04
+SR 82.75 ± 0.03 92.31 ± 0.07 91.87 ± 0.03 91.45 ± 0.05
+RI 75.89 ± 0.05 93.19 ± 0.12 92.35 ± 0.05 93.88 ± 0.13
+RS 92.13 ± 0.06 94.11 ± 0.08 91.67 ± 0.14 90.11 ± 0.06
+RD 90.15 ± 0.08 94.67 ± 0.09 93.03 ± 0.02 92.89 ± 0.09
+BT 85.62 ± 0.04 90.12 ± 0.13 92.55 ± 0.09 91.12 ± 0.03
+MixCode 92.62 ± 0.01 95.91 ± 0.01 92.58 ± 0.06 -
+StarCoder2-7B GenCode 92.91 ± 0.03 96.32 ± 0.12 93.03 ± 0.04 94.21 ± 0.12
+No Aug 92.61 ± 0.03 94.35 ± 0.11 92.82 ± 0.14 93.45 ± 0.06
+SR 85.67 ± 0.04 92.67 ± 0.13 91.93 ± 0.08 91.55 ± 0.08
+RI 78.94 ± 0.06 93.01 ± 0.21 92.65 ± 0.06 93.75 ± 0.04
+RS 91.59 ± 0.07 94.31 ± 0.09 92.04 ± 0.02 90.32 ± 0.09
+RD 90.65 ± 0.11 95.01 ± 0.04 92.65 ± 0.07 93.14 ± 0.12
+BT 83.65 ± 0.02 91.23 ± 0.11 92.86 ± 0.13 91.56 ± 0.05
+MixCode 92.89 ± 0.09 94.89 ± 0.02 92.67 ± 0.03 -
+Qwen2.5-Coder-7B GenCode 93.12 ± 0.02 96.17 ± 0.15 92.93 ± 0.09 94.38 ± 0.14
+Maximum probability analysis.
+To explore why GenCode produces a more robust code model, we check the confidence of trained models using different code augmentation methods on the original test data.
+Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data Wu et al.
+(2018).
+To do so, for each dataset and code model, we collect the maximum probability of the corrected classified test data produced by the code model first.
+Then, we calculate the mean values of these maximum probability scores.
+The results of the mean values of maximum probabilities are shown in Table 8.
+We can see that code models trained by GenCode are more confident in the test data than models trained using other code augmentation methods.
+As a result, GenCode trained code models are harder to attack.
 Table 8: Mean values of maximum probabilities.
@@
 CodeBERT GenCode 0.0413779 0.9808122 0.9714346 0.9786421
+Statistical analysis.
+We also conduct statistical tests using Wilcoxon signed-rank test.
+When comparing existing data augmentation methods, GenCode consistently exhibits a p-value << 0.05 across all pre-trained code models shown from the lower of Table 5.
+In conclusion, GenCode can produce more adversarial robust code understanding models than other methods.
+Answer to RQ2: Our statistical tests find GenCode can train more robust pre-trained code models than existing data augmentation methods.
+On average, code models trained using GenCode have 8.42% and 4.90% higher robustness than models trained without data augmentation and models trained using MixCode, respectively.
+However, for LLMs, the robustness improvement achieved by GenCode remains limited, with gains of less than 1%.
+7.3 RQ3: How does the influence score affect the effectiveness of GenCode?
+Even though we have preliminarily shown that selecting the code samples with higher loss values is a valid way to train good models, it is still necessary to study to what extent the advantage can be brought from this selection strategy.
+Therefore, in this part, we conduct an ablation study to compare GenCode with its two variants, GenCode with data selection using minimum loss values - GenCode (Min) and random selection - GenCode (Random), respectively.
 Table 9: Effectiveness of data augmentation methods w.r.t.
-test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data and ASR $\downarrow$ (%) on test data.
+test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data.
+The best results are highlighted in blue.
+Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
+Model DA method GCJ Refactory Java250 BigCloneBench
+Random 94.74 ± 0.21 98.82 ± 0.15 98.24 ± 0.11 98.27 ± 0.19
+Min 93.34 ± 0.32 96.13 ± 0.21 96.43 ± 0.17 97.23 ± 0.14
+CodeBERT Max 96.23 ± 0.18 99.32 ± 0.24 98.89 ± 0.09 98.93 ± 0.15
+Random 95.26 ± 0.11 98.93 ± 0.18 98.68 ± 0.22 98.67 ± 0.13
+Min 93.53 ± 0.17 96.31 ± 0.24 96.72 ± 0.12 97.22 ± 0.21
+GraphCodeBERT Max 97.93 ± 0.14 99.45 ± 0.17 99.16 ± 0.13 99.23 ± 0.09
+Random 96.24 ± 0.21 98.99 ± 0.21 98.93 ± 0.16 98.83 ± 0.08
+Min 94.11 ± 0.31 98.29 ± 0.12 97.03 ± 0.23 97.01 ± 0.15
+CodeT5 Max 98.85 ± 0.16 99.64 ± 0.37 99.38 ± 0.11 99.34 ± 0.12
+Random 96.81 ± 0.06 99.11 ± 0.02 99.23 ± 0.04 99.31 ± 0.06
+Min 95.42 ± 0.03 98.68 ± 0.04 98.47 ± 0.03 98.13 ± 0.03
+StarCoder2-7B Max 98.93 ± 0.04 99.65 ± 0.02 99.41 ± 0.05 99.36 ± 0.09
+Random 97.12 ± 0.05 99.12 ± 0.03 99.23 ± 0.03 99.32 ± 0.04
+Min 95.86 ± 0.08 98.71 ± 0.01 98.51 ± 0.05 98.19 ± 0.07
+Qwen2.5-Coder-7B Max 98.96 ± 0.04 99.65 ± 0.01 99.42 ± 0.04 99.36 ± 0.05
+Table 9 and Table 10 summarize the accuracy and robustness of trained code models using different selection strategies.
+We can see GenCode consistently outperforms the two variants under all testing criteria.
+Selecting code samples with minimum loss values is the worst way for GenCode.
+Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on the results.
+Table 10: Effectiveness of data augmentation methods w.r.t.
+ASR $\downarrow$ (%) on test data.
 The best results are highlighted in blue.
@@
 GCJ Refactory Java250 BigCloneBench
-Model DA method Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$
-GenCode (Random) 94.74 ± 0.21 25.21 50.32 98.82 ± 0.15 27.65 36.65 98.24 ± 0.11 32.52 44.87 98.27 ± 0.19 8.01 18.12
-GenCode (Min) 93.34 ± 0.32 35.66 66.67 96.13 ± 0.21 65.51 69.76 96.43 ± 0.17 46.76 53.86 97.23 ± 0.14 20.65 41.32
-CodeBERT GenCode 96.23 ± 0.18 22.33 42.62 99.32 ± 0.24 22.75 29.25 98.89 ± 0.09 29.89 40.65 98.93 ± 0.15 7.12 15.83
-GenCode (Random) 95.26 ± 0.11 29.97 50.63 98.93 ± 0.18 17.56 24.29 98.68 ± 0.22 20.65 38.86 98.67 ± 0.13 2.04 5.97
-GenCode (Min) 93.53 ± 0.17 40.55 68.23 96.31 ± 0.24 38.73 46.98 96.72 ± 0.12 26.87 46.67 97.22 ± 0.21 11.45 21.23
-GraphCodeBERT GenCode 97.93 ± 0.14 26.43 39.87 99.45 ± 0.17 15.78 20.92 99.16 ± 0.13 18.76 33.81 99.23 ± 0.09 1.86 4.62
-GenCode (Random) 96.24 ± 0.21 25.03 47.61 98.99 ± 0.21 18.46 23.67 98.93 ± 0.16 25.11 42.34 98.83 ± 0.08 8.67 17.42
-GenCode (Min) 94.11 ± 0.31 38.65 59.89 98.29 ± 0.12 35.67 48.59 97.03 ± 0.23 32.61 49.27 97.01 ± 0.15 18.36 25.54
-CodeT5 GenCode 98.85 ± 0.16 23.57 38.76 99.64 ± 0.37 12.13 18.81 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
-Visualizing principal component analysis.
-Furthermore, we explore why models trained by GenCode have higher accuracy than models trained using other methods by analyzing the code representation capability of trained code embeddings.
-Specifically, 1) first, we extract the code embeddings produced by each trained model using all the test data as code features.
-2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) (Maćkiewicz and Ratajczak 1993) algorithm.
-3) After that, we visualize the reduced 2-dimensional vectors.
-Normally, the features extracted by a well-trained model are more clearly distributed based on the data labels.
-Figure 3 depicts the visualized features extracted from CodeBERT using the Refactory dataset.
-We can see that the code embeddings trained by using GenCode are more clearly distributed than embedding extracted by other models, i.e., the buggy code and correct code are greatly different in the figure.
-These results indicate that GenCode can train good code embeddings and, therefore, produce better accurate code models for code understanding tasks.
-Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Figure.
-3.
-The results from Table 4 confirm that GenCode can better distinguish the two classes.
-Interestingly, we found that MixCode has the smallest distance result.
-That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries.
-This phenomenon is similar to the findings in (Verma et al.
-2019), where interpolating hidden state representations yields two beneficial properties: 1) class representations are flattened into a minimal number of directions of variation, and 2) all points between these flat representations are assigned low-confidence predictions.
-Robustness Evaluation
-Overall robustness.
-Column MHM and ALERT in Table 3 summarize the attack success rate of adversarial attack methods against code models.
-First, we can see that similar to the observation in Accuracy Evaluation, not all data augmentation methods can improve the robustness of code models.
-This finding is consistent with the existing works (Dong et al.
-2023b).
-Specifically, in the baseline methods, only RS and MixCode have a positive impact on the robustness of code models.
-Considering our method, GenCode consistently improves the robustness of code models and achieves the best results in 21 out of 24 cases.
-On average, compared to No Aug, models trained by GenCode have an 8.42% lower ASR (higher robustness).
-This result breaks the previous finding (Bielik and Vechev 2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness.
-On the other hand, compared to MixCode, GenCode can train code models with 4.90% higher robustness on average.
-Statistical analysis.
-We also conduct statistical tests using T-Test and Wilcoxon signed-rank test.
-When comparing existing data augmentation methods, GenCode consistently exhibits a p-value $<$ 0.05 across all pre-trained models.
-In conclusion, GenCode can produce more adversarial robust code understanding models than other methods.
-Maximum probability analysis.
-To explore why GenCode produces a more robust code model, we check the confidence of trained models using different code data augmentation methods on the original test data.
-Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data (Wu et al.
-2018).
-To do so, for each dataset and code model, we collect the maximum probability of the corrected classified test data produced by the code model first.
-Then, we calculate the mean values of these maximum probability scores.
-The results of the mean values of maximum probabilities are shown in Table 5.
-We can see that code models trained by GenCode are more confident in the test data than models trained using other code data augmentation methods.
-As a result, GenCode trained code models are harder to attack.
-Ablation Study
-Even though we have preliminarily shown that selecting the code samples with higher loss values is a valid way to train good models, it is still necessary to study to what extent the advantage can be brought from this selection strategy.
-Therefore, in this part, we conduct an ablation study to compare GenCode with its two variants, GenCode with data selection using minimum loss values - GenCode (Min) and random selection - GenCode (Random), respectively.
-Table 6 summarizes the accuracy and robustness of trained code models using different selection strategies.
-We can see GenCode consistently outperforms the two variants under all testing criteria.
-Selecting code samples with minimum loss values is the worst way for GenCode.
-Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on the results.
-This phenomenon indicates simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods is already a good code data augmentation method.
-Discussion
+Model DA method MHM ALERT MHM ALERT MHM ALERT MHM ALERT
+Random 25.21 50.32 27.65 36.65 32.52 44.87 8.01 18.12
+Min 35.66 66.67 65.51 69.76 46.76 53.86 20.65 41.32
+CodeBERT Max 22.33 42.62 22.75 29.25 29.89 40.65 7.12 15.83
+Random 29.97 50.63 17.56 24.29 20.65 38.86 2.04 5.97
+Min 40.55 68.23 38.73 46.98 26.87 46.67 11.45 21.23
+GraphCodeBERT Max 26.43 39.87 15.78 20.92 18.76 33.81 1.86 4.62
+Random 25.03 47.61 18.46 23.67 25.11 42.34 8.67 17.42
+Min 38.65 59.89 35.67 48.59 32.61 49.27 18.36 25.54
+CodeT5 Max 23.57 38.76 12.13 18.81 21.63 39.76 7.39 12.15
+This phenomenon indicates that simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods, is already a good code ugmentation method.
+Answer to RQ3: Selecting code samples with maximum loss values is the best strategy among our considered variants, i.e., with minimum loss values and random selection.
+Besides, GenCode with random selection can outperform existing code augmentation methods.
+7.4 RQ4: How does the data selection method affect the effectiveness of GenCode?
+Table 11 presents the results of accuracy.
+Overall, LIME demonstrates limited effectiveness and generally underperforms compared to the gradient-based method (i.e.,  BADGE).
+Influence Function achieves better performance in scenarios involving smaller datasets, such as CodeT5 on GCJ and GraphCodeBERT on Refactory.
+Additionally, BADGE exhibits superior performance on LLMs, particularly in tasks such as Qwen2.5-Coder-7B on Refactory and BigCloneBench.
+Although GenCode outperforms other methods in 12 out of 20 cases, these results suggest that exploring combinations of existing data selection strategies could be a promising direction for future research.
+Table 11: Effectiveness of data augmentation methods w.r.t.
+test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data.
+The best results are highlighted in blue.
+Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
+Model Data Selection GCJ Refactory Java250 BigCloneBench
+LIME 93.67 ± 0.21 98.81 ± 0.13 97.25 ± 0.19 97.45 ± 0.14
+Influence Function 95.65 ± 0.19 99.38 ± 0.09 97.13 ± 0.24 97.58 ± 0.19
+BADGE 96.15 ± 0.13 99.34 ± 0.22 98.21 ± 0.17 98.95 ± 0.12
+CodeBERT GenCode 96.23 ± 0.18 99.32 ± 0.24 98.89 ± 0.09 98.93 ± 0.15
+LIME 94.98 ± 0.07 98.12 ± 0.17 98.01 ± 0.09 97.83 ± 0.09
+Influence Function 97.98 ± 0.11 99.46 ± 0.02 98.79 ± 0.08 97.79 ± 0.13
+BADGE 96.88 ± 0.07 98.65 ± 0.06 99.03 ± 0.11 98.41 ± 0.04
+GraphCodeBERT GenCode 97.93 ± 0.14 99.45 ± 0.17 99.16 ± 0.13 99.23 ± 0.09
+LIME 96.36 ± 0.11 98.45 ± 0.06 99.03 ± 0.04 98.21 ± 0.13
+Influence Function 98.97 ± 0.09 99.23 ± 0.13 99.29 ± 0.13 99.31 ± 0.12
+BADGE 98.86 ± 0.18 99.54 ± 0.16 99.36 ± 0.03 99.33 ± 0.18
+CodeT5 GenCode 98.85 ± 0.16 99.64 ± 0.37 99.38 ± 0.11 99.34 ± 0.12
+LIME 96.96 ± 0.08 99.13 ± 0.04 99.25 ± 0.09 99.32 ± 0.04
+Influence Function 97.35 ± 0.05 99.41 ± 0.06 99.31 ± 0.05 99.34 ± 0.03
+BADGE 98.96 ± 0.14 99.65 ± 0.04 99.39 ± 0.12 99.35 ± 0.06
+StarCoder2-7B GenCode 98.93 ± 0.04 99.65 ± 0.02 99.41 ± 0.05 99.36 ± 0.09
+LIME 97.46 ± 0.05 99.25 ± 0.12 99.04 ± 0.02 99.32 ± 0.13
+Influence Function 98.13 ± 0.07 99.52 ± 0.05 99.12 ± 0.07 99.35 ± 0.02
+BADGE 98.98 ± 0.06 99.49 ± 0.03 99.38 ± 0.02 99.37 ± 0.03
+Qwen2.5-Coder-7B GenCode 98.96 ± 0.04 99.65 ± 0.01 99.42 ± 0.04 99.36 ± 0.05
+Table 12: Effectiveness of data augmentation methods w.r.t.
+test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on transformed test data.
+The best results are highlighted in blue.
+Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
+Model Data Selection GCJ Refactory Java250 BigCloneBench
+LIME 89.13 ± 0.07 95.41 ± 0.11 93.21 ± 0.14 91.68 ± 0.12
+Influence Function 90.34 ± 0.05 95.51 ± 0.07 93.35 ± 0.12 92.17 ± 0.15
+BADGE 91.31 ± 0.09 95.61 ± 0.05 93.61 ± 0.03 92.64 ± 0.11
+CodeBERT GenCode 91.23 ± 0.14 95.63 ± 0.14 93.62 ± 0.09 92.56 ± 0.15
+LIME 88.24 ± 0.06 95.35 ± 0.07 94.24 ± 0.02 93.62 ± 0.08
+Influence Function 88.97 ± 0.08 95.62 ± 0.03 94.32 ± 0.07 93.58 ± 0.12
+BADGE 90.01 ± 0.04 95.88 ± 0.05 94.87 ± 0.11 93.96 ± 0.07
+GraphCodeBERT GenCode 89.78 ± 0.16 95.67 ± 0.15 94.89 ± 0.08 93.89 ± 0.11
+LIME 90.45 ± 0.13 95.85 ± 0.13 94.32 ± 0.04 93.43 ± 0.05
+Influence Function 91.04 ± 0.16 95.94 ± 0.11 94.68 ± 0.09 93.65 ± 0.09
+BADGE 91.35 ± 0.14 96.14 ± 0.15 95.14 ± 0.09 93.95 ± 0.14
+CodeT5 GenCode 91.38 ± 0.09 95.87 ± 0.11 95.06 ± 0.12 94.03 ± 0.12
+LIME 92.72 ± 0.16 94.56 ± 0.14 92.67 ± 0.12 93.91 ± 0.09
+Influence Function 92.89 ± 0.03 95.13 ± 0.19 92.86 ± 0.11 94.12 ± 0.15
+BADGE 92.93 ± 0.04 96.24 ± 0.17 93.12 ± 0.13 94.19 ± 0.19
+StarCoder2-7B GenCode 92.91 ± 0.03 96.32 ± 0.12 93.03 ± 0.04 94.21 ± 0.12
+LIME 92.94 ± 0.07 95.13 ± 0.04 92.68 ± 0.12 93.87 ± 0.05
+Influence Function 92.88 ± 0.19 95.78 ± 0.09 92.76 ± 0.08 94.19 ± 0.06
+BADGE 93.09 ± 0.05 96.19 ± 0.02 92.97 ± 0.11 94.32 ± 0.05
+Qwen2.5-Coder-7B GenCode 93.12 ± 0.02 96.17 ± 0.15 92.93 ± 0.09 94.38 ± 0.14
+An interesting observation from the natural robustness results in Table 12 is that, in most cases, the gradient-based method achieves better robustness improvements than GenCode.
+Specifically, BADGE outperforms GenCode by integrating gradient-based uncertainty with diversity via K-means clustering, thereby facilitating the selection of informative and representative samples Ash et al.
+(2019).
+Answer to RQ4: GenCode achieves better improvements in accuracy compared to other data selection methods.
+However, the gradient-based method demonstrates superior performance in enhancing robustness.
+Investigating the trade-offs between GenCode and gradient-based methods presents an interesting direction for future research.
+8 Discussion
+8.1 Computational cost of augmented data
+We conduct a case study to evaluate the computational cost of model training with different data augmentation methods.
+The results are visualized in Fig.
+6.
+We observe that GenCode incurs the highest computational cost compared to other data augmentation methods including code refactoring, text-oriented, and MixCode.
+Interestingly, text-oriented data augmentation methods consistently exhibit lower training-time cost than code refactoring methods.
+In particular, Random Deletion (RD) achieves the best trade-off between performance and efficiency.
+Although Back Translation (BT) performs relatively well across some tasks (e.g., authorship attribution and bug detection), it requires significant preprocessing time to generate the augmented data before model training, making it less practical.
+Therefore, we do not recommend BT as a preferred augmentation strategy in time-sensitive scenarios.
+(a) Authorship attribution
+(b) Bug detection
+(c) Problem classification
+(d) Clone detection
+Figure 6: Computational cost of CodeBERT using different code augmentation methods in each task.
+In conclusion, rule-based augmentation methods (e.g., code refactoring or text-oriented data augmentation methods) typically require less computation during model training, but may incur the preprocessing overhead.
+In contrast, sample-based methods such as MixCode do not require data preparation in advance and can be flexibly applied during training.
+While GenCode consistently achieves the best accuracy and robustness, it comes with the highest computational cost.
+Exploring the trade-off between performance and efficiency is an important direction for future work.
+8.2 On the Data Distribution of Augmented Data
+One common concern in data augmentation is the risk of introducing samples that deviate from the original data distribution, potentially harming model generalization.
+While augmented data may not strictly follow the original distribution, prior studies (e.g., Mixup Zhang et al.
+(2018), CutMix Yun et al.
+(2019)) have shown that moderate, controlled deviations can actually enhance generalization.
+To mitigate distributional shift, GenCode follows a principled two-stage pipeline: (1) controlled data generation, and (2) loss-based importance-guided selection.
+In the first stage, GenCode applies a combination of semantic-preserving code refactorings (e.g., variable renaming, field enhancement) and lightweight text-oriented data augmentation methods(e.g., random swap, synonym replacement).
+These transformations are designed to preserve or minimally alter program semantics, ensuring that the generated data remains close to the original distribution.
+In the second stage, GenCode evaluates each augmented sample based on its training loss under the current model.
+Only samples with the highest loss, those that are potentially more informative, are selected for training, while others are discarded.
+This loss-based filtering mechanism promotes alignment with the model’s effective learning region and helps avoid training on overly noisy or unrepresentative data.
+8.3 Limitations
 Even though the accuracy of our studied code models already achieved relatively high accuracy in these tasks, i.e., most of them have more than 90% accuracy, their robustness is still low.
-Regarding the large language models (LLMs), we have checked the literature and found that the clone detection task is still challenging for LLMs, e.g., ChatGPT has only 0.784 Precision for Java-Java pair detection (Khajezade et al.
-2024).
+Regarding the large language models (LLMs), we have checked the literature and found that the clone detection task is still challenging for LLMs, e.g., ChatGPT has only 0.784 Precision for Java-Java pair detection Khajezade et al.
+(2024).
 For other program understanding tasks that have a large number of classes, e.g., the Java250 dataset used in our experiments has 250 classes.
-It is also hard to use open-source LLMs such as ChatGPT to predict the label since we need to add many label guidance in the prompt.
-Conclusion
+It is also hard to use LLMs such as ChatGPT to predict the label since we need to add many label guides in the prompt.
+8.4 Threats to Validity
+The internal threat to validity mainly comes from the implementation of standard model training and data augmentation methods.
+The code that is used in our experiment includes 1) the authorship attribution task (Dataset: GCJ) taken from the existing work Yang et al.
+(2022), 2) the bug detection task (Dataset: Refactory) adopted from Hu et al.
+(2019), 3) the program classification task (Dataset: Java250) sourced from Project CodeNeet Puri et al.
+(2021), 4) the clone detection task (Dataset: BigCloneBench) chosen from the work  Svajlenko et al.
+(2014).
+The implementation of SR, RI, RS, RD, and BT source from the newest release Marivate and Sefara (2020), again we adapt them to our experiment.
+The implementation of MixCode and Refactor come from its original release Dong et al.
+(2023).
+The external threats to validity in our study are associated with the code understanding tasks, datasets, pre-trained code models, and data augmentation methods we have selected.
+We have considered four distinct code understanding tasks: authorship attribution, bug detection, problem classification, and clone detection, utilizing a total of four datasets for these tasks.
+Our study encompasses two widely used programming languages in the software community, namely Java and Python.
+Furthermore, we employ three types of pre-trained code models, namely CodeBERT, GraphCodeBERT, and CodeT5, as well as two recently released code-specific large language models, StarCoder2 and Qwen2.5-Coder, in our experiments.
+For data augmentation methods from NLP, we have selected established techniques based on their widespread citation in academic papers and their recognition in authoritative journals and conferences.
+The construct threats to validity in our study primarily come from various factors, including parameter settings, randomness, and the choice of evaluation measures.
+For MixCode, the primary parameter is lambda, which governs the weight of mixing two feature vectors.
+We have adhered to the original recommendation for this parameter.
+Similarly, for data augmentation methods from NLP, we have used the parameters as per their original release.
+To mitigate the influence of randomness, we have conducted each experiment five times and reported the results as an average along with the standard deviation.
+In terms of evaluation metrics, our study encompasses both accuracy and robustness.
+While accuracy assesses the basic performance of pre-trained code models, robustness is employed to evaluate their generalization capabilities.
+9 Related Work
+Data augmentation has made remarkable success in the field of machine learning, e.g., natural language processing Feng et al.
+(2021) and code model Zhuo et al.
+(2023).
+Indeed, data augmentation has proven to be a valuable technique for improving model performance, specifically increasing trained model accuracy and enhancing dataset diversity Rebuffi et al.
+(2021).
+Inspired by that code can be represented as text data Allamanis et al.
+(2018), we review related works from three perspectives, deep learning for code understanding, data augmentation for code models, and data augmentation for NLP.
+9.1 Deep Learning for Code Understanding
+Code understanding is important for software developers during the development and maintenance phases.
+Benefiting from the potential of using deep learning models to understand natural language texts, researchers from the SE borrow the same idea to build deep learning models for code understanding.
+Code2vec Alon et al.
+(2019b) is the very first work that introduced the use of DNN models to learn vector information extracted by code snippets.
+Code2vec achieved promising results in understanding code on the function name prediction task.
+Later on, CodeBERT Feng et al.
+(2020), was proposed for more code-understanding tasks.
+CodeBERT follows the pre-training and fine-tuning paradigm.
+That means it uses multiple programming languages to pre-train the model to learn generation code information.
+Then, people can fine-tune the pre-trained CodeBERT on a specific downstream task for real usage.
+CodeBERT achieved impressive results on many code understanding such as comments generation.
+After that, multiple pre-trained code models have been proposed with ever better performance, such as GraphCodeBERT Guo et al.
+(2020), GraphCode2Vec Ma et al.
+(2022), and CodeT5 Wang et al.
+(2021).
+More recently, large language models such as ChatGPT  [^7] https://chat.openai.com/ have been demonstrated to be the best choice to solve code understanding tasks.
+Different from the above works that focused on designing new model architectures to understand code.
+We target a more general problem – how to prepare the training data to further boost the training of code models.
+Our approach is suitable for any code understanding model.
+9.2 Data Augmentation for code models
+Inspired by the success of data augmentation, researchers have recently dedicated significant efforts to applying data augmentation techniques to large-scale code-related tasks to enhance the performance of code models, focusing on accuracy and robustness.
+Adversarial training Goodfellow et al.
+(2015), which involves generating a set of adversarial examples and incorporating them into the training data, has been explored as a data augmentation method in code learning.
+To enhance the performance of deep comment generation models, Zhang et al.
+Zhang et al.
+(2021) proposed a code augmentation method that utilizes the Metropolis-Hastings Modifier (MHM) algorithm to generate the adversarial training dataset.
+Similar to their work, Mi et al.
+Mi et al.
+(2021) generated additional data using Auxiliary Classifier Generative Adversarial Networks (GANs).
+Besides, as a method that is tailored specifically for code-related tasks, code refactoring which involves restructuring and optimizing code without changing its external behavior is becoming a mainstream code augmentation.
+Yu et al.
+Yu et al.
+(2022) devised program transformation rules tailored for Java and conducted an extensive assessment of their effectiveness when employed as code augmentation techniques across three significant code-related tasks.
+Allamanis et al.
+Allamanis et al.
+(2021) applied four simple code rewrite rules to improve the efficiency of model training in the program repair task.
+Furthermore, taking inspiration from Mixup, which linearly mixes the features of two distinct image datasets, Dong et al.
+Dong et al.
+(2023) proposed Mixcode that linearly interpolates the features derived from a pair of programs along with their corresponding labels, wherein One of the programs undergoes code refactoring transformation, while the other remains unchanged.
+Li et al.
+Li et al.
+(2022) also introduced alternative interpolation methods for the code model, namely binary interpolation and linear extrapolation.
+Unlike previous work, our study is the first one that considers the importance of code data for enhancing the performance of code understanding tasks.
+Specifically, during each epoch, the most valuable data selected by the metric score are fed into the code model for enhancing the performance of training.
+9.3 Data Augmentation for NLP
+Data augmentation methods that are used in NLP can be roughly divided into three categories, namely paraphrasing, noising-based methods, and sampling-based methods Feng et al.
+(2021).
+Xie et al.
+Xie et al.
+(2020) proposed the back-translation technique that involves bidirectionally applying a translation model to each sentence.
+More precisely, this method first translates the original text into another language and then reverts it back to the original one.
+Paraphrasing is widely employed in the field of NLP since it can convey identical information as the original text but in a different form.
+Wei et al.
+Wei and Zou (2019) applied a series of nosing-based methods named EDA to augment the text data.
+Noising-based methods generally involve the subtle addition of noise to the original data while preserving their semantic information.
+For example, Random Swap (one of the methods from EDA) can include randomly selecting two words in a sentence and swapping their positions.
+Besides, different from paraphrasing or noising-based methods, sampling-based methods are task-specific and necessitate both data and their label formats.
+Also inspired by the Mixup, Guo et al.
+Guo et al.
+(2019) adapted it from its original use with image data to text data.
+They conducted an empirical study to investigate its effectiveness.
+Specifically, they generated new synthetic data by linearly mixing the embeddings instead of directly operating on the raw text data.
+Furthermore, to overcome the limitations of traditional methods such as Synonym Replacement, which generally require significant prior knowledge and are prone to sub-optimal results, Ren et al.
+Ren et al.
+(2021) proposed Text AutoAugment.
+In detail, it considers a combination of various text-oriented data augmentation operations as an augmentation policy and employs an efficient Bayesian Optimization algorithm to automatically search for the best policy.
+In contrast to the data augmentation methods used in NLP, our work focuses on the strategy that guides code model training.
+The strategy involves both existing data augmentation methods derived directly from code data and those adapted from text data.
+Moreover, our approach also allows for the integration of new program transformation methods that researchers may develop in the future.
+10 Conclusion
 We propose the first generation-and-selection-based framework, GenCode, to boost code understanding models.
-The basic assumption behind GenCode is that the less confident data contributes more to the model training.
+The basic assumption behind GenCode is that less confident data contributes more to the model training.
 Based on this assumption, GenCode utilizes atomic data augmentation operators to generate candidate training data first and selects data with maximum loss values to train the code model.
-Our evaluation demonstrates that GenCode can produce more accurate and robust models than MixCode with 2.92% accuracy and 4.90% robustness improvement, respectively.
-Besides, we analyze how the importance score affects the effectiveness of GenCode.
-In the future, we plan to 1) explore the combination of GenCode with other confidence scores, such as uncertainty metrics and active learning methods; 2) study the effectiveness of GenCode on fine-tuning open-source large language models, e.g., Llama.
+Our evaluation demonstrates that GenCode can produce more accurate and robust models than MixCode.
+Besides, we analyze how the influence score affects the effectiveness of GenCode.